### v1

# An Optimism-based Approach to Online Evaluation of Generative Models

Abstract
Existing frameworks for evaluating and comparing generative models typically target an offline setting, where the evaluator has access to full batches of data produced by the models. However, in many practical scenarios, the goal is to identify the best model using the fewest generated samples to minimize the costs of querying data from the models. Such an online comparison is challenging with current offline assessment methods. In this work, we propose an online evaluation framework to find the generative model that maximizes a standard assessment score among a group of available models. Our method uses an optimism-based multi-armed bandit framework to identify the model producing data with the highest evaluation score, quantifying the quality and diversity of generated data. Specifically, we study the online assessment of generative models based on the Fréchet Inception Distance (FID) and Inception Score (IS) metrics and propose the FID-UCB and IS-UCB algorithms leveraging the upper confidence bound approach in online learning. We prove sub-linear regret bounds for these algorithms and present numerical results on standard image datasets, demonstrating their effectiveness in identifying the score-maximizing generative model.
1 Introduction
Deep generative models have achieved astonishing results across a wide array of machine learning datasets. Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) [17] and Inception Score (IS) [37]. Due to the growing applications of generative models to various learning tasks, the machine learning community has continuously adapted evaluation methodologies to better suit the characteristics of the newly introduced applications.
A common characteristic of standard evaluation frameworks for deep generative models is their offline assessment process, which requires a full batch of generated data for assigning scores to the models. While this offline evaluation does not incur significant costs for moderate-sized generative models, producing large batches of samples from large-scale models can be costly. In particular, generating a large batch of high-resolution image or video data could be expensive and hinder the application of existing evaluation scores for ranking generative models.
In this work, we focus on the online evaluation of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data. By limiting the number of generated samples in the assessment process, online evaluation can save on the costs associated with producing large batches of samples. Additionally, online evaluation can significantly reduce the time and computational expenses required to identify a well-performing model in assessing a large group of generative models.
To measure the performance of an online evaluation algorithm, we use the regret notion widely-used in the online learning literature. In the online learning task, the evaluation algorithm selects a generative model and observes a mini-batch of samples generated by that model in each round. Our goal is to minimize the algorithm’s regret, defined as the cumulative difference between the scores of the selected models and the best possible score from the model set. Therefore, the learner’s regret quantifies the cost of generating samples from sub-optimal models. Our goal is to develop online learning algorithms that result in lower regret values in the assessment of generated data.
To address the described online evaluation of generative models, we utilize the multi-armed bandit framework. In a standard multi-armed bandit problem, the online learner seeks to identify the arm with the highest expected value of a random score. While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FID metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
By deriving concentration bounds for the FID and IS scores, we propose the optimism-based FID-UCB and IS-UCB algorithms to address the online assessment of generative models. These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FID and IS scores. Furthermore, we analytically bound the regret of the IS-UCB and FID-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
We discuss the results of several numerical applications of FID-UCB and IS-UCB to standard image datasets and generative modeling frameworks. We compare the performance of these algorithms with the greedy algorithm, which selects the model with the highest estimated score, and a naive-UCB baseline that applies the UCB algorithm with a data-independent upper confidence bound. Our numerical results show a significant improvement in our proposed algorithms compared to the greedy and naive-UCB baselines. Additionally, we analyze the performance of FID-UCB with different image data embeddings, including InceptionNet.V3 [40], CLIP [7], and DINOv2 [33]. Our empirical results indicate satisfactory performance under these standard embeddings. The following summarizes the main contributions of this work:
  • Proposing an online evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
  • Developing the FID-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
  • Proving sub-linear regret bounds for the proposed FID-UCB and IS-UCB algorithms.
  • Demonstrating satisfactory empirical performance of FID-UCB and IS-UCB in comparison to the greedy and naive-UCB baseline algorithms.
2 Related Work
Assessment of deep generative models. The evaluation of generative models has been extensively studied in the literature. Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic [3], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [5], and diversity/quality-based metrics such as Precision/Recall [36, 27], density and coverage [32], VENDI [13], and RKE [20]. In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score [2], the FLD score [21], the Rarity score [16], and the KEN score [45] measuring the novelty of the generated samples. In this paper, we primarily focus on FID and Inception score, which have been frequently used for evaluating generative models.
Role of embeddings in the quantitative evaluation results. Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 [40]. However, [32] shows that such pretrained-embeddings could exhibit unexpected behaviors. Recently, several large pretrained models have been proposed, including DINOv2 [33] and CLIP [7]. [39] shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models. In addition, [28] demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments. In this paper, we provide the numerical results for the mentioned embeddings extracted by different pretrained models.
Online Learning using diversity-related evaluation metrics. Online learning is a sequential decision-making framework where an agent aims to minimize a cumulative loss function revealed to her sequentially. One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of [29, 41]. At each step, the agent chooses among several arms, each associated with a reward distribution, and aims to maximize a pre-specified performance metric. The primary concern of this body of literature considers maximizing the expected return [1, 4, 6]. Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion [38, 47] in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards [44]. However, to the best of our knowledge, the evaluation of generative models has not been exclusively studied in an online learning context.
Online training of generative models. Several related works focus on training generative adversarial networks (GANs) [14] using online learning frameworks. [15] proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach. [10] shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs). Also, the recent paper [34] studies the no-regret behaviors of large language model (LLM) agents. This reference proposes an unsupervised training loss, whose minimization could automatically result in known no-regret learning algorithms. On the other hand, our focus is on the evaluation of generative models which does not concern the models’ training.
3 Preliminaries
3.1 Inception Score
Inception score (IS) is a standard metric for evaluating generative models, defined as
$\textup{IS}:=\exp\left\{\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[\textup{KL}(p_{Y|X_{\textup{g}}}\|p_{Y_{\textup{g}}})]\right\},$ (1)
where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet [40], and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution. Further, we have that $\textup{IS}=\exp[I(X_{\textup{g}};Y_{\textup{g}})]=\exp[H(Y_{\textup{g}})-H(Y_{\textup{g}}|X_{\textup{g}})]$, where $I(\cdot;\cdot)$ is the mutual information, and $H(\cdot)$ is the Shannon entropy. A higher IS implies that the images generated by the model have higher diversity, since $p_{Y_{\textup{g}}}$ would be more uniformly distributed to increase $H(Y_{\textup{g}})$, and possess higher fidelity, because $p_{Y|X_{\textup{g}}}$ is closer to a one-hot vector to enforce a smaller $H(Y_{\textup{g}}|X_{\textup{g}})$.
3.2 Fréchet Inception Distance
Fréchet Inception Distance (FID) is another standard metric for evaluating generative models. Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$, which is extracted by the InceptionNet.V3 embedding. FID is the Fréchet distance [12] between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, defined as
$\textup{FID}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}]$ (2)
assuming that $p_{f(X_{\textup{g}})}={\mathcal{N}}(\mu_{\textup{g}},\Sigma_{\textup{g}})$ and $p_{f(X_{\textup{r}})}={\mathcal{N}}(\mu_{\textup{r}},\Sigma_{\textup{r}})$ are Gaussian distributions.
4 Online Evaluation of Generative Models
In this section, we introduce the framework of online evaluation of generative models, which is given in Protocol 1. We denote by ${\mathcal{G}}:=[G]$ the set of generative models. For each generator $g\in[G]$, we denote by $p_{g}\in\Delta({\mathcal{X}})$ its generative distribution over the space ${\mathcal{X}}$, which can be texts or images. Given an evaluation metric, e.g., FID or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$. The evaluation proceeds in $T$ steps. At each step $t\in[T]$, the evaluating algorithm ${\mathcal{A}}$ picks a generator $g_{t}\in[G]$ and collects a batch of generated samples $\{x_{t}^{j}\sim p_{g}\}_{j=1}^{b}$, where $b\in{\mathbb{N}}_{+}$ is the (fixed) batch size. The evaluator aims to minimize the regret
$\textup{Regret}(T)=\sum_{t=1}^{T}(\nu^{*}-\nu_{g_{t}}),$ (3)
where $\nu^{*}:=\operatorname*{arg\,max}_{g\in[G]}\nu_{g}$ (if the higher the score the better).
Regarding the challenges of online evaluation of generative models, note that the empirical estimation of the score could be biased and generator-dependent. In addition, the generative distribution typically lies in a high-dimensional space, which makes it difficult to estimate the score from limited data. Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FID requires the covariance matrix). Hence, analyzing the concentration properties of the score estimation would be challenging.
Protocol 1 Online Evaluation of Generative Models
[…]
1:  Initialize: estimated score $\widehat{\nu}_{g}$ and generated samples ${\mathcal{H}}_{g}\leftarrow\varnothing$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$  do
3:     Pick generator $g_{t}\sim{\mathcal{A}}$.
4:     Generate a batch $\{x_{t}^{j}\sim p_{g_{t}}\}_{j=1}^{b}$ of samples from $g_{t}$.
5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{x_{t}^{j}\}_{j=1}^{b}$ and estimated score $\widehat{\nu}_{g_{t}}$.
6:  end for
5 Online Evaluation based on Fréchet Inception Distance
In this section, we consider the online evaluation of generative models by Fréchet Inception Distance (FID). Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FID is computed by
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)
where
$\widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top},\qquad\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i})$ (5)
are the empirical covariance matrix and the mean vector, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3. It has been shown that the empirical FID (4) is biased differently depending on the generator, which makes it difficult to compare different generators for a fixed sample size [9]. Hence, the FID-based evaluation is typically performed for a large batch of generated samples (e.g., $10\sim 50$k) to reduce the effect of the bias, which can be sample-inefficient and costly for a large number of generators.
To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FID score. To this end, we first derive an optimistic FID score in the following theorem. We defer the theorem’s proof to Appendix A.1.
Theorem 5.1 (Generator-dependent optimistic FID score).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive definite. Then, with probability at least $1-\delta$, we have
$\widehat{\textup{FID}}_{g}^{n}=\widetilde{\textup{FID}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FID}_{g}.$ (6)
Here, the bonus is given by
${\mathcal{B}}_{g}^{n}:=$ $C^{n}_{g}\cdot\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (7)
$+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\frac{2{\mathcal{I}}_{g}}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)},$
where $C_{g}^{n}:=2(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})$, ${\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$, and $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$.
Algorithm 1 FID-UCB for Generative Model Evaluation
[…]
1:  Initialize: $\widehat{\textup{FID}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$ do
3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FID}}_{g}$.
4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\in{\mathbb{R}}^{D}\sim g_{t}\}_{j=1}^{b}$.
5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
6:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
7:     Compute the optimistic FID score
$\widehat{\textup{FID}}_{g_{t}}\leftarrow\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}.$
8:  end for
We remark that the bonus term (7) is: 1) generator-dependent through parameters including ${\mathcal{I}}_{g}$, $\sigma_{g}^{2}$, $\textup{Tr}[\Sigma_{g}]$, and $\|\Sigma_{g}\|_{2}$, 2) dimension-free in the sense that the dimension $d$ only appears in the logarithmic term, which facilitates a sample-efficient evaluation. In practice, these parameters can be estimated from the queried data.
Based on Theorem 5.1, we propose FID-UCB in Algorithm 1 as an FID-based online evaluation algorithm. At the beginning, the estimated FID scores are initialized to be negative infinity (line 1), and hence each generator is explored for at least one time. The evaluation proceeds iterative, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FID and queries a batch of images from the model (lines $3\sim 4$). Then, the estimated FID of generator $g_{t}$ is updated (lines $5\sim 7$). Next, we show that FID-UCB attains sub-linear regret bound. We defer the theorem’s proof to Appendix A.2.
Theorem 5.2 (Regret of FID-UCB).
Under the same conditions in Theorem A.1, with probability at least $1-\delta$, the regret of the UCB-FID algorithm after $T$ steps is bounded by
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$ (8)
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
6 Online Evaluation based on Inception Score
In this section, we focus on evaluating generative models by Inception score (IS), which is given by $\textup{IS}=\exp[I(X_{g};Y_{g})]$, where $X_{g}\sim p_{g}$ is the generated image, and $Y_{g}$ is the class assigned by the InceptionNet.V3. Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, the empirical IS is computed by
$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\},$ (9)
where
$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),\ \widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (10)
are the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$ ($d=1000$ for InceptionNet.V3).
For any $j\in[d]$, let $\widehat{V}^{n}(p_{Y|X_{g}}[j]):={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ denote the sample variance for the probability that the generated sample is assigned to the $j$-th class. To derive an optimistic IS, we first define the optimistic marginal class distribution denoted by
$\widehat{p}_{Y_{g}}^{n}:=\textup{Clip}_{e^{-1}}(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g})$ (11)
where $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ is a $d$-dimensional vector whose $j$-th element is given by
$\bm{\epsilon}^{n}_{g}[j]:=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$ (12)
and $\textup{Clip}_{e^{-1}}(p,\bm{\epsilon})$ is the following element-wise operator
$\begin{cases}p[j]+\frac{e^{-1}-p[j]}{|e^{-1}-p[j]|}\bm{\epsilon}[j]&\textup{, if }|e^{-1}-p[j]|\geq\bm{\epsilon}[j]\\ e^{-1}&\textup{, otherwise.}\end{cases}$ (13)
The optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n}):=-\sum_{j}\widehat{p}_{Y_{g}}^{n}[j]\cdot\log(\widehat{p}_{Y_{g}}^{n}[j])\geq H(Y_{g})$ with high probability (see Lemma B.1 in the Appendix B.3). Next, we derive a generator-dependent optimistic IS in the following theorem. We defer the theorem’s proof to Appendix B.1.
Theorem 6.1 (Generator-dependent Optimistic IS).
Let
$\widehat{\textup{IS}}^{n}_{g}:=\exp\left\{{\mathcal{E}}\left(\widehat{p}_{Y_{g}}^{n}\right)-\widehat{H}^{n}(Y_{g}|X_{g})+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)\right\},$ (14)
where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (10), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (11), parameter $C^{\prime}$ satisfies that $P(|H(Y_{g}|X_{g})|\leq C^{\prime})=1$, and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution. Then, with probability at least $1-\delta$, we have that $\widehat{\textup{IS}}^{n}_{g}\geq\textup{IS}_{g}$.
[…]
1:  Initialize: $\widehat{\textup{IS}}_{g}\leftarrow\infty,\widehat{H}(Y_{g}|X_{g})\leftarrow 0,\widetilde{p}_{Y_{g}}\leftarrow\bm{0}^{d},n_{g}\leftarrow 0$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$ do
3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{G}}}\widehat{\textup{IS}}_{g}$.
4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\sim p_{g_{t}}\}_{j=1}^{b}$.
5:     Update
$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},\ \widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
6:     Update visitation $N_{g_{t}}\leftarrow N_{g_{t}}+b$.
7:     Update the empirical variances $\widehat{V}(H(Y_{g_{t}}|X_{g_{t}}))$ and $\widehat{V}(p_{Y|X_{g_{t}}}[j])$ for any $j\in[d]$.
8:     Compute the estimated IS according to Equation (14).
9:  end for
Algorithm 2 IS-UCB for Generative Model Evaluation
Based on Theorem 6.1, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm. In the following theorem, we derive the regret bound of the IS-UCB algorithm, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret. We present the theorem’s proof in Appendix B.2.
Theorem 6.2 (Regret of IS-UCB).
With probability at least $1-\delta$, the regret of the IS-UCB algorithm after $T$ steps is bounded by
$\textup{Regret}(T)\leq\widetilde{O}\left(e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}+dG\Delta^{-2}\log d\right),$ (15)
where $C:=O(de^{-1}+C^{\prime})$ is the upper bound of $\widehat{\textup{IS}}_{g}^{n}$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
7 Experimental Results
Baseline methods. For both FID-based and IS-based evaluation, we compare the performance of our proposed FID-UCB and IS-UCB with the following two baselines, naive-UCB and the greedy algorithm. Naive-UCB is a simplification of FID-UCB and IS-UCB which is based on data-independent dimension-based concentration bounds, which do not exploit the generated data in evaluating the confidence bound. We present the exact estimators in Appendix D. On the other hand, the greedy algorithm always picks the generator with the lowest empirical FID (4) or the highest empirical IS (9). For all these algorithms, we initialize the estimated FID to be $\infty$- ∞ and the estimated IS to be +$\infty$+ ∞, and hence each generator will be explored at least once.
[…]
Experimental setup and performance metrics. We report three performance metrics at each step $t\in[T]$: 1) total regret, i.e., $\textup{Regret}(t)$, 2) average regret, i.e., $(1/t)\cdot\textup{Regret}(t)$, and 3) optimal pick ratio, i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FID score or the highest empirical IS score for 50k generated images. For all the experiments, we use a batch size of 5, and the total evaluation step is $T=1,000$. Hence the total generated samples for each trial is $5$k. Results are averaged over 20 trials. Due to the limited space, we postpone parts of the results to Appendix C.
7.1 Empirical Results of Online FID-based Evaluation
Performance on pretrained generators under different embeddings. The results are summarized in Figure 1, where each row and each column correspond to one dataset and one performance metric, respectively. The results show that FID-UCB (blue) outperforms naive-UCB (orange) and the greedy algorithm (green). We observe that naive-UCB fails to distinguish between different generators and makes uniform selections on the CIFAR10 and the ImageNet datasets, while FID-UCB recognizes the optimal one. This shows that the generator-dependent optimistic FID (6) can better exploit the properties of the generator, which is key to the sample-efficient online evaluation. On the FFHQ dataset, all the generators attain very similar FID scores (the largest gap is less than 0.7), which may explain why FID-UCB has a similar performance to naive-UCB. We reran these experiments using DINOv2-ViT-L/14-based [33] and CLIP-based [7] embeddings. (The results are summarized in Figures 5 and 6 in Appendix C, respective.) The numerical results show that FID-UCB can maintain satisfactory performance over different embeddings.
CIFAR10 (G=5)
ImageNet (G=5)
[…]
Figure 1: Performance of FID-UCB, naive-UCB, and the greedy algorithm on various datasets. The embeddings of the images are extracted by the pretrained InceptionNet.V3.
Performance on variance-controlled generators. We summarize the results in Figure 2. FID-UCB (blue) significantly outperforms naive-UCB and the greedy algorithm. The results for DINOv2-based and CLIP-based FID are provided in Figures 5 and 6 in Appendix C. We ran the evaluation on AFHQ Dog dataset, and the results are presented in Figure 7 in the Appendix. Furthermore, we display the generated images to visualize the quality and diversity of the samples queried by different algorithms. The results suggest that FID-UCB consistently queries images that are diverse and of high quality. In contrast, naive-UCB kept querying a proportion of images from the collapsed generator (the one with the smallest truncation parameter), and the greedy algorithm even stuck to it.
FFHQ (G=20)
Figure 2: Performance of FID-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
Step 500
[…]
7.2 Results of Online IS-Based Evaluation
Performance on the pretrained generators. The results are summarized in Figure 3. IS-UCB (blue) attains significantly better performance than naive-UCB (orange) and the greedy algorithm (green) on CIFAR10 and FFHQ datasets.
CIFAR10 (G=5)
Figure 3: Performance of IS-UCB, naive-UCB, and the greedy algorithm on various datasets.
ImageNet (G=5) [Uncaptioned image]
[…]
Performance on variance-controlled generators. The results are summarized in Figure 4. Generated samples are plotted in Figure 8 in the Appendix. The results on the AFHQ Dog dataset are presented in Figure 9 in the Appendix.
FFHQ (G=20)
Figure 4: Performance of IS-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
8 Conclusion
In this work, we studied an online learning problem aiming to identify the generative model with the best evaluation score among a set of models. We proposed the optimism-based FID-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models. An interesting future direction to our work is to explore the application of other online learning frameworks, such as Thompson sampling, to address the online evaluation problem. Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FID-UCB and IS-UCB to text-based and video-based generative models will be interesting for future studies. Such an application only requires applying text and video-based embeddings, as in Video IS [35], FVD [42], and FTD [19].
Appendix A Proofs in Section 5: FID Evaluation
A.1 Proof of Theorem 5.1: Optimistic FID Score
Proof.
The proof of Theorem 5.1 is based on Theorem A.1 and Lemma A.2 in the Appendix. Specifically, Inequality (16) in Theorem A.1 in Appendix A.3 decomposes the estimation error of the empirical FID score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$). Next, Lemma A.2 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix. Combining ∎
A.2 Proof of Theorem 5.2: Regret of FID-UCB
Proof.
Recall that $g_{t}$ denotes the generator picked at the $t$-th step. For convenience, we denote by $\widehat{\textup{FID}}_{g_{t}}$ the optimistic FID of generator $g_{t}$ computed at step $t$. We denote by $n_{t}$ the number of images generated by model $g_{t}$ after $t$ steps. First, by Theorem A.2 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FID}_{g}\geq\widehat{\textup{FID}}_{g}$ for any step $t\in[T]$ and $g\in[G]$. Hence, we have that
$\textup{Regret}(T)=\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\textup{FID}^{*}\right)\leq\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widehat{\textup{FID}}_{g_{t}}\right).$
By the definition of $\widehat{\textup{FID}}_{g_{t}}$, we further derive that
$\textup{Regret}(T)\leq$ $\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widetilde{\textup{FID}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)\leq 2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}},$
where ${\mathcal{B}}^{n}_{g}$ is the bonus function given by Equation (7). Therefore, the regret is further bounded by
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])\cdot T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Here, we use the fact that
$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{(2b)^{\alpha}}{1-\alpha}\sum_{g=1}^{G}(n_{g}(T))^{1-\alpha}\leq\frac{(2b)^{\alpha}}{1-\alpha}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq 2b\sum_{g=1}^{G}\log(n_{g}(T))\leq 2bG\log(T),$
where $b$ is the batch size, and $n_{g}(T)$ is the number of images generated by model $g$ at the last step $T$. ∎
A.3 Concentration of Empirical FID (4)
Theorem A.1 (Concentration of empirical FID (4)).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite, and we denote by $\Sigma_{\textup{r}}=Q_{\textup{r}}^{\top}\Lambda_{\textup{r}}Q_{\textup{r}}$ its eigendecomposition. Then, with probability at least $1-\frac{\delta}{3}$, we have that
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$ (16)
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}.$
Proof.
Recall that for any generator $g$, we have that
$\textup{FID}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  Note that
$\left|\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}-\|\mu_{g}-\mu_{\textup{r}}\|_{2}\right|\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
We derive
$\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}\right|$ (17)
$\leq$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
$\leq$ $(2\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.  By Theorem A.3 in Appendix A.5, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}\left[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (18)
(3) Bound $\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]$.  Note that
$\left|\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]\right|$
$=$ $\left|\mathbb{E}[\textup{Tr}[(f(X_{g})-\mu_{g})(f(X_{g})-\mu_{g})^{\top}]]-\frac{1}{n}\sum_{i=1}^{n}\textup{Tr}[(f(x^{i})-\widehat{\mu}_{g}^{n})((f(x^{i})-\widehat{\mu}_{g}^{n}))^{\top}]\right|$
$=$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2}\right|$
$\leq$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|.$ (19)
Next, we bound the two terms in the last inequality. Let $Z_{g}:=f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. For the first term, we have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\|Z_{g}\|_{2}^{2}=(\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Sigma_{g}(\Sigma_{g}^{-\frac{1}{2}}Z_{g})=(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Lambda_{g}(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}),$
where we denote by $\Sigma_{g}:=Q_{g}^{\top}\Lambda_{g}Q_{g}$ the eigendecomposition of $\Sigma_{g}$. Note that the random vector $Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}$ has distribution ${\mathcal{N}}(0,I)$. We have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\sum_{j=1}^{d}\lambda_{j}U_{j}^{2},$
where $\lambda_{j}$’s are the eigenvalues of $\Sigma_{g}$, $U_{i}\sim{\mathcal{N}}(0,1)$ are independent standard normal, and hence $U_{i}^{2}$’s are independent Chi-square random variables. Therefore, we have that
$\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|$
$=$ $\left|\sum_{j=1}^{d}\lambda_{j}\left(\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-\mathbb{E}[U_{j}^{2}]\right)\right|$
$\leq$ $\left(\sum_{j=1}^{d}\lambda_{j}\right)\cdot\max_{j=1,\cdots,d}\left|\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-1\right|$
$\leq$ $\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)},$ (20)
where the last inequality holds with probability at least $1-\frac{\delta}{3}$ by Lemma E.5. In addition, for the second term, we derive that
$\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|$
$=$ $\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})(\|f(x^{i})-\mu_{g}\|_{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right|$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right)$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right).$ (21)
Putting all together.  Therefore, combining Inequalities (17), (18), and (19), with probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}},$
which concludes the proof. ∎
A.4 Concentration of Mean Vector and Covariance Matrix
Lemma A.2 (Concentration of mean vector and covariance matrix).
Under the same conditions in Theorem A.1, with probability at least $1-\frac{2\delta}{3}$, we have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}},\text{ where }{\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$ (22)
and
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2},$ (23)
where $\sigma_{g}^{2}:=\max_{i\in[d]}\Sigma_{g}[i,i]$ is the largest variance for the elements in $f(X_{g})$, and $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
Proof.
1. Concentration of mean vector.  Note that $\Sigma_{g}^{\frac{1}{2}}Z$ where $Z\sim{\mathcal{N}}(0,I)$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution, which is the same as $f(X_{g})-\mu_{g}$. We have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}(\Sigma_{g}^{\frac{1}{2}}Z_{i})\right\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}Z_{i}\right\|_{\Sigma_{g}},$
where $\|y\|_{\Sigma_{g}}:=\sqrt{\langle y,\Sigma_{g}y\rangle}$ is the matrix norm. Let $\widehat{Z}:=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ denote the mean vector. We further obtain that
$\left\|\widehat{Z}\right\|_{\Sigma_{g}}^{2}=\textup{Tr}\left((\widehat{Z})^{\top}\Sigma_{g}\widehat{Z}\right)=\textup{Tr}\left(\Sigma_{g}\widehat{Z}(\widehat{Z})^{\top}\right)=\textup{Tr}(\Sigma_{g}D)+\textup{Tr}(\Sigma_{g}M),$ (24)
where $D:=\textup{diag}(\widehat{Z}(\widehat{Z})^{\top})$ and $M=\widehat{Z}(\widehat{Z})^{\top}-D$. For the first term, note that each diagonal element has the same distribution $\frac{1}{n^{2}}(\sum_{i=1}^{n}z_{i})^{2}$, where $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\frac{\delta}{6}$, we have that
$\textup{Tr}(\Sigma_{g}D)\leq\|D\|_{2}\textup{Tr}(\Sigma_{g})\leq\textup{Tr}(\Sigma_{g})\left(\frac{1}{n}\sqrt{32\log\left(\frac{12d}{\delta}\right)}\right),$ (25)
where the first inequality holds by that $\Sigma_{g}$ is PSD, and the second inequality holds by Lemma E.10. For the second term,
$\textup{Tr}(\Sigma_{g}M)=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot M_{j,l}$
$=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)$
$\leq$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\cdot\max_{j,l}\left|\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)\right|$
where $z_{j,1},\cdots,z_{j,n},z_{l,1},\cdots,z_{l,n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. By Lemma E.11, with probability at least $1-\frac{\delta}{6}$, it holds that
$\textup{Tr}(\Sigma_{g}M)\leq\left(\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\right)\frac{1}{n}\sqrt{8\log\left(\frac{24d}{\delta}\right)}$ (26)
Combining Inequalities (25) and (26), we conclude that with probability at least $1-\frac{\delta}{3}$, it holds that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|^{2}_{2}\leq\left(\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|\right)\left(\frac{2}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)$ (27)
2. Concentration of sample covariance matrix. we have that
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}$
$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}$
$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}$
$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2},$ (28)
where the last inequality holds by the triangle inequality of the L2 norm of the matrix. Note that the zero-mean random vector $f(X_{g})-\mu_{g}$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution (under the assumption of FID). For the first term, we evoke Lemma E.7. Then, with probability at least $1-\frac{\delta}{3}$, it holds that
$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}.$ (29)
For the second term, we have that
$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2},$
which concludes the proof. ∎
A.5 Dimension-free Concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
Theorem A.3 (Dimension-free concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
Under the same conditions in Theorem A.1, we have that
$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (30)
Proof.
Let $\eta=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$. We have that
$\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B}+\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
$\leq$ $\textup{Tr}\left[\sqrt{2\eta B^{\top}B}\right]=\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\eta},$
where the first inequality holds by Lemma E.10, the second inequality holds by Lemma E.11 and the fact that both matrices $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B\succeq\bm{0},B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by Lemma E.10 and $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$. By the same analysis, we can also derive that
$\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]\leq\textup{Tr}\left[\sqrt{2\eta B^{\top}B}\right]=\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\eta},$
which concludes the proof. ∎
Appendix B Proofs in Section 6: IS Evaluation
B.1 Proof of Theorem 6.1: Generator-Dependent Optimistic IS
Proof.
By Lemma B.2, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 6.1, it suffices to show that
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}-\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)$
with probability at least $1-\frac{\delta}{2}$, which holds by Theorem E.4. Therefore, we conclude the proof. ∎
B.2 Proof of Theorem 6.2: Regret of IS-UCB
Proof.
By Theorem 6.1 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g}\geq\textup{IS}_{g}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (14). For convenience, we denote by $\widehat{\textup{IS}}_{g_{t}}$ the optimistic IS of generator $g_{t}$ computed at the $t$-th step. Hence, we have that
$\textup{IS}^{*}-\textup{IS}_{g_{t}}\leq\widehat{\textup{IS}}_{g_{t}}-\textup{IS}_{g_{t}}=e^{{\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}}-\textup{IS}_{g_{t}},$
where
${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right).$
Let ${\mathcal{T}}_{1},{\mathcal{T}}_{2}\subset[T]$ be two disjoint set of steps such that ${\mathcal{T}}_{2}=[T]\backslash{\mathcal{T}}_{1}$. Specifically, ${\mathcal{T}}_{1}$ contains steps where no element of the empirical marginal class distribution $\widetilde{p}_{Y_{g_{t}}}$ is clipped to $e^{-1}$, i.e., $|e^{-1}-\widetilde{p}_{Y_{g_{t}}}[j]|\geq\bm{\epsilon}_{g_{t}}[j]$ for all $[j]\in d$. Let $\textup{IS}^{*}=\operatorname*{arg\,max}_{g\in[G]}\textup{IS}_{g}$ denote the optimal Inception score. Hence, we have that
$\textup{Regret}(T)=\sum_{t\in{\mathcal{T}}_{1}}(\textup{IS}^{*}-\textup{IS}_{g_{t}})+\sum_{t\in{\mathcal{T}}_{2}}(\textup{IS}^{*}-\textup{IS}_{g_{t}}).$
Recall $u(x)=-x\log x$ for $x\in{\mathbb{R}}_{+}$. For the first part, we further derive that
$\sum_{t\in{\mathcal{T}}_{1}}(\textup{IS}^{*}-\textup{IS}_{g_{t}})$
$\leq$ $\sum_{t\in{\mathcal{T}}_{1}}\left(\exp\{{\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}\}-\exp\{H(Y_{g})-H(Y_{g}|X_{g})\}\right)$
$\leq$ $e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left({\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}-H(Y_{g})+H(Y_{g}|X_{g})\right)$
$\leq$ $e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(\sum_{j=1}^{d}\left(u(\widehat{p}_{Y_{g_{t}}}[j])-u(p_{Y_{g_{t}}}[j])\right)+2{\mathcal{B}}_{g_{t}}\right)$
$\leq$ $O\left(e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(m\cdot\sum_{j=1}^{d}|\widehat{p}_{Y_{g_{t}}}[j]-p_{Y_{g_{t}}}[j]|+{\mathcal{B}}_{g_{t}}\right)\right)$
$\leq$ $O\left(e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(m\cdot\sum_{j=1}^{d}\bm{\epsilon}_{g_{t}}[j]+{\mathcal{B}}_{g_{t}}\right)\right)$
$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left(\max\left\{dm,\frac{d\log d}{2}\right\}\sqrt{\frac{1}{N_{g_{t}}}}+\frac{C^{\prime}}{N_{g_{t}}}\right)\right),$ (31)
where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=O(de^{-1}+C^{\prime})$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$. For the second part, note that the number of steps where at least one element of $\widetilde{p}_{Y_{g_{t}}}$ is clipped to at most $\widetilde{O}(G\Delta^{-2})$, where $\Delta=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$. Therefore, we have that
$\textup{Regret}(T)\leq\widetilde{O}\left(dG\Delta^{-2}\log d+e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}\right),$
which concludes the proof. ∎
B.3 Optimistic marginal class distribution
Theorem B.1 (Optimistic marginal class distribution).
Let $\widetilde{p}^{n}_{Y_{g}}:=\frac{1}{n}\sum_{i=1}^{n}p_{Y_{g}|x^{i}}$ denote the empirical marginal class distribution. Let
$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},|\widetilde{p}^{n}_{Y_{g}}-p_{Y_{g}}|\right),$
where $\textup{Clip}_{e^{-1}}(p,\bm{\epsilon})$ is the following element-wise operator
$\begin{cases}p[j]+\frac{e^{-1}-p[j]}{|e^{-1}-p[j]|}\bm{\epsilon}[j]&\textup{, if }|e^{-1}-p[j]|\geq\bm{\epsilon}[j]\\ e^{-1}&\textup{, otherwise.}\end{cases}$
for any vectors $p,\bm{\epsilon}\in{\mathbb{R}}^{d}_{+}$. Then, we have that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
Note that the function $u(z)=-z\log z$ is concave and attains maximum at $z=e^{-1}$. It suffices to show that $u(\widehat{p}_{Y_{g}}[j])\geq u(p_{Y_{g}}[j])$ for all $j\in[d]$. If $|e^{-1}-\widehat{p}_{Y_{g}}^{n}[j]|<|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|$, then $\widehat{p}_{Y_{g}}[j]=e^{-1}$, which ensures that $u(\widehat{p}_{Y_{g}}[j])\geq u(p_{Y_{g}}[j])$. In addition, if $|e^{-1}-\widehat{p}_{Y_{g}}^{n}[j]|\geq|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|$, then
$\widehat{p}_{Y_{g}}^{n}[j]=\widetilde{p}^{n}_{Y_{g}}[j]+\frac{e^{-1}-\widetilde{p}^{n}_{Y_{g}}[j]}{|e^{-1}-\widetilde{p}^{n}_{Y_{g}}[j]|}|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|=\begin{cases}\widetilde{p}^{n}_{Y_{g}}[j]+|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|,&\textup{ if }p_{Y_{g}}[j]<e^{-1}\\ \widetilde{p}^{n}_{Y_{g}}[j]-|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|,&\textup{ otherwise.}\end{cases}.$
The first case ensures that $p_{Y_{g}}[j]<\widehat{p}_{Y_{g}}^{n}[j]\leq e^{-1}$, and the second case ensures that $p_{Y_{g}}[j]>\widehat{p}_{Y_{g}}^{n}[j]\geq e^{-1}$. Both cases satisfy that $u(\widehat{p}_{Y_{g}}^{n}[j])\geq u(p_{Y_{g}}[j])$. Therefore, we have that $H(Y_{g})={\mathcal{E}}(p_{Y_{g}})=\sum_{j=1}^{d}u(p_{Y_{g}}[j])\leq\sum_{j=1}^{d}u(\widehat{p}_{Y_{g}}^{n}[j])={\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})$, which concludes the proof. ∎
B.4 Data-dependent optimistic marginal class distribution
Lemma B.2 (Optimistic marginal class distribution).
Let $x^{1},\cdots,x^{n}\sim p_{g}$ be $n$ generated images from generator $g$. Define $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ denote the error vector whose $j$-th element is given by
$\bm{\epsilon}^{n}_{g}[j]=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
where $\widehat{V}^{n}(p_{Y|X_{g}}[j])={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ is the empirical variance for the $j$-th class density. Then, with probability at least $1-\frac{\delta}{2}$, we have that
$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g}\right)$
satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ with probability at least $1-\frac{\delta}{2}$. We evoke Theorem E.4 and conclude the proof. ∎
Appendix C Additional Experimental Results
C.1 Results of FID-Based Online Evaluation
1. Embeddings Extracted by DINOv2 [33].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by DINOv2-ViT-L/14. The pretrained weights are downloaded from the official repository (licensed under Apache License 2.0). [^3] https://github.com/facebookresearch/dinov2/tree/main The results are summarized in Figure 5.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 5: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of the images are extracted by the pretrained DINOv2-ViT-L/14 model.
2. Embeddings Extracted by CLIP [7].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by CLIP. The pretrained weights are downloaded from the official repository. [^4] https://github.com/mlfoundations/open_clip. Copyright (c) 2012-2021 Gabriel Ilharco, Mitchell Wortsman, Nicholas Carlini, Rohan Taori, Achal Dave, Vaishaal Shankar, John Miller, Hongseok Namkoong, Hannaneh Hajishirzi, Ali Farhadi, Ludwig Schmidt. The results are summarized in Figure 6. We note that in the FFHQ dataset, FID-UCB attains a low optimal pick ratio as the optimal generator and the second-best one attains almost the same CLIP-based FID.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 6: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of images are extracted by the pretrained CLIP model.
3. Results on AFHQ Dog dataset.
The results are summarized in Figure 7.
AFHQ Dog (G=20)
Figure 7: Performance of FID-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ Dog dataset. The embeddings of images are extracted by InceptionNet.V3.
C.2 Results of IS-Based Online Evaluation
1. Diversity and quality of the generated samples.
We examine the diversity and quality of the generated samples queried by IS-UCB, naive UCB, and the greedy algorithm (i.e., experiment in Figure 4). The results are summarized in Figure 8.
Step 500
[…]
Figure 8: Generated images queried by IS-UCB, naive UCB, and the greedy algorithm.
2. Results on AFHQ Dog dataset.
The results are summarized in Figure 9.
AFHQ Dog (G=20)
Figure 9: Performance of IS-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ dataset.
Appendix D Additional Experimental details
1. Additional experiment setups.
All the experiments are run on a single Nvidia GeForce RTX 3090 GPU. For FID-based evaluation, each trial takes at most 2 hours to finish, depending on the dataset and the embedding pretrained model. For IS-based evaluation, each trial takes less than 30 minutes to finish.
2. List of pretrained generative models/generated image data.
For the CIFAR10 dataset, we compare pretrained models including SNGAN, ACGAN, BigGAN, ContraGAN, and BigGAN-CR. For the ImageNet dataset, we compare pretrained models including StyleGAN3, SNGAN, BigGAN, ContraGAN, and ReACGAN. The pretrained weights are downloaded from the StudioGAN repository [22]. For the FFHQ dataset, we compare StyleSwin, StyleNAT, StyleGAN2-ADA, StyleGAN-XL, and LDM, where we utilize the generated image datasets downloaded from the dgm-eval repository [39].
3. Implementation details of naive UCB for FID-based evaluation.
In FID-based evaluation, naive UCB upper bounds the error in estimating the mean vector by
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\sqrt{\sum_{j=1}^{d}(\widehat{\mu}_{g}^{n}[j]-\mu_{g}[j])^{2}}\leq\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{2d}{\delta}\right)}$
which holds by the assumption of FID score that $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$, with probability at least $1-\delta$. In addition, the error in estimating the covariance matrix is bounded by
$\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}+\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}\right)^{2}\right)$
which follows from the exact same derivation in Inequality (28) and holds with probability at least $1-\delta$ (see Lemma E.6). Compared to the generator-dependent error bounds utilized by FID-UCB (i.e., Inequalities (27) and (29)), these bounds are dimension-dependent and do not fully exploit the properties of the generator. Finally, by the exact same analysis in the proof of Theorem 5.1, the following estimated FID
$\widehat{\textup{FID}}_{g}^{n}=$ $\widetilde{\textup{FID}}_{g}^{n}-C_{g}^{n}\cdot\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{6d}{\delta}\right)}-\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (32)
$-\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}\right)^{2}\right)+\frac{d\sigma_{g}^{2}}{n}\log\left(\frac{6d}{\delta}\right)\right)}$
lower bounds $\textup{FID}_{g}$ with probability at least $1-\delta$.
4. Implementation details of naive UCB for IS-based evaluation.
In IS-based evaluation, naive UCB uses Hoeffding-type bonus terms. Specifically, the error vector in defining the optimistic marginal class distribution is given by $e_{g}^{n}$, whose $j$-th element is defined as
$e_{g}^{n}[j]:=\sqrt{\frac{1}{2n}\log\left(\frac{4d}{\delta}\right)},$
which ensures that ${\mathcal{E}}(\textup{Clip}_{e^{-1}}(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}))\geq H(Y_{g})$ with probability at least $1-\frac{\delta}{2}$. In addition, the concentration error of $\widehat{H}^{n}(Y_{g}|X_{g})$ is bounded by
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}$
which holds with probability at least $1-\frac{\delta}{2}$. Therefore, the estimated IS is given by
$\widehat{\textup{IS}}_{g}^{n}:=\exp\left\{{\mathcal{E}}\left(\textup{Clip}_{e^{-1}}\left(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}\right)\right)-\widehat{H}^{n}(Y_{g}|X_{g})+C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}\right\}$ (33)
By the exact same analysis in the proof of Theorem 6.1, it can be shown that $\widehat{\textup{IS}}_{g}^{n}\geq\textup{IS}_{g}$ with probability at least $1-\delta$.
5. Additional implementation details of FID-UCB
To compute the bonus term (7), we need to estimate the following parameters: 1) $\|\mu_{g}-\mu_{\textup{r}}\|_{2}$: we set it to be a constant, e.g., 0.5; 2) ${\mathcal{I}}_{g}$: we estimate this term using the truncated covariance matrix with parameter $\tau$, which is denoted by $\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})$. The $(i,j)$-th element of the truncated covariance matrix is $\widehat{\Sigma}_{g}^{n}[i,j]$ if $\widehat{\Sigma}_{g}^{n}[i,j]\geq\tau$ and is zero otherwise. In the experiments, we observe that a truncation parameter that is $5\sim 10\%$ of the operator norm $\|\Sigma_{\textup{r}}\|_{2}$, i.e., $0.05\cdot\|\Sigma_{\textup{r}}\|_{2}\leq\tau\leq 0.1\cdot\|\Sigma_{\textup{r}}\|_{2}$, attains satisfying performance; 3) $\|\Sigma_{g}\|_{2}$: we estimate this term by $\max_{j\in[d]}(\widehat{\sigma}_{j}^{n})^{2}+O(\sqrt{\log(1/\delta)/n})$, where $(\widehat{\sigma}_{j}^{n})^{2}$ is the sample variance for the $j$-th entry of the embedding $f(X_{g})$. This avoids the computation for the eigendecomposition of the sample covariance matrix $\widehat{\Sigma}_{g}^{n}$. The intuition is that when $\Sigma_{g}$ is diagonal, the operator norm $\|\Sigma_{g}\|_{2}$ coincides with the largest per-entry variance; 4) $\textup{Tr}[\Sigma_{g}]$: while a direct approach is to estimate it by $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$, we found that a better performance can be attained by using $O(\textup{Tr}[\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})\cdot\Sigma_{\textup{r}}])$, where we utilize the fact that $\|\Sigma_{\textup{r}}^{-1}\|_{2}\cdot\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]\geq\textup{Tr}[\Sigma_{g}]\geq\frac{\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]}{\|\Sigma_{\textup{r}}\|_{2}}$.
Appendix E Auxiliary Lemmas
Lemma E.1 (Concentration of Gaussian random variable).
Let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|>\epsilon\right)<\exp\left(-\frac{n\epsilon^{2}}{2\sigma^{2}}\right)$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|\leq\sqrt{\frac{\sigma^{2}}{2n}\log\left(\frac{2}{\delta}\right)}.$
Lemma E.2 (Hoeffding’s inequality).
Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d. random variables with values in $[0,1]$ and let $\delta>0$. Then with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{1}{2n}\log\left(\frac{1}{\delta}\right)}.$ (34)
A drawback of Hoeffding’s inequality is that the confidence interval is distribution-agnostic, and always scales with order $O(\sqrt{1/n})$. For distribution with small variance, the following [18, Bennett’s inequality] provides a tighter concentration error.
Lemma E.3 (Bennett’s inequality).
Under the conditions of Theorem E.2, with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2{\mathbb{V}}Z}{n}\log\left(\frac{1}{\delta}\right)}+\frac{1}{3n}\log\left(\frac{1}{\delta}\right),$ (35)
where ${\mathbb{V}}Z$ is the variance ${\mathbb{V}}=\mathbb{E}(Z-\mathbb{E}Z)^{2}$.
When the exact variance is unknown, [31] provide a purely data-dependent bound with similar properties as Bennett’s inequality.
Lemma E.4 (Restatement of [31, Theorem 4]).
Under the conditions of Theorem E.2, we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2V_{n}(\bm{Z})}{n}\log\left(\frac{2}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{2}{\delta}\right),$
where $V_{n}(\bm{Z}):=\frac{1}{n(n-1)}\sum_{1\leq i<j\leq n}(Z_{i}-Z_{j})^{2}$ is the sample variance.
Lemma E.5 (Concentration of Chi-square variables).
Let $Z_{1},\cdots,Z_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard normal. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|>\epsilon\right)\leq 2e^{-n\epsilon^{2}/8}.$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|\leq\sqrt{\frac{8}{n}\log\left(\frac{2}{\delta}\right)}$
Lemma E.6 (Restatement of [43, Example 6.3]).
Let $\bm{X}\in{\mathbb{R}}^{n\times d}$ be a random matrix from the $\Sigma$-Gaussian ensemble, i.e., each row vector $x_{i}$ is drawn i.i.d. from a multivariate ${\mathcal{N}}(0,\Sigma)$ distribution. Then, with probability at least $1-\delta$, we have that
$\|\widehat{\Sigma}-\Sigma\|_{2}\leq\|\Sigma\|_{2}\left(2\sqrt{\frac{d}{n}}+2\epsilon+\left(\sqrt{\frac{d}{n}}+\epsilon\right)^{2}\right),\textup{ where }\epsilon=\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)},$ (36)
where $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\top}$ is the sample covariance matrix.
Lemma E.7 (Dimension-free concentration of sample covariance matrix [30, 25, 46]).
Let $X_{1},\cdots,X_{n}$ are i.i.d zero-mean and sub-Gaussian random vectors with parameter at most $\sigma^{2}$, whose covariance matrix is given by $\Sigma$. Let
$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (37)
denote the effective rank. Then, with probability at least $1-\delta$, the sample covariance matrix satisfies that
$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\sigma^{2}\|\Sigma\|_{2}\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}$ (38)
whenever $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$.
Lemma E.8.
Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
$\textup{Tr}(\sqrt{A+B})\geq\max\{\textup{Tr}(\sqrt{A}),\textup{Tr}(\sqrt{B})\}$ (39)
Proof.
Since $A+B\succeq A$ and $A+B\succeq B$, by Courant-Fischer min-max theorem, it holds that $\lambda_{i}(A+B)\geq\lambda_{i}(A)$ and $\lambda_{i}(A+B)\geq\lambda_{i}(B)$ for their $i$-th largest eigenvalues, which concludes the proof. ∎
Lemma E.9.
Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
$\textup{Tr}[\sqrt{A+B}]\leq\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}].$ (40)
Proof.
First, we assume that both matrices $A$ and $B$ are positive strictly definite. Then,
$\textup{Tr}(\sqrt{A+B})=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}(A+B)(A+B)^{-\frac{1}{4}}]$
$=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}A(A+B)^{-\frac{1}{4}}]+\textup{Tr}[(A+B)^{-\frac{1}{4}}B(A+B)^{-\frac{1}{4}}]$
$=$ $\textup{Tr}[\sqrt{A}(A+B)^{-\frac{1}{2}}\sqrt{A}]+\textup{Tr}[\sqrt{B}(A+B)^{-\frac{1}{2}}\sqrt{B}]$
$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma E.12). For general PSD matrices $A$ and $B$, note that
$\textup{Tr}(\sqrt{A+B+\epsilon I})\leq\textup{Tr}(\sqrt{A+\epsilon I})+\textup{Tr}(\sqrt{B+\epsilon I})$
holds for any $\epsilon>0$ by our previous analysis. Since the trace map is continuous, ∎
Lemma E.10.
Let $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}\leq\sqrt{\frac{32}{n^{2}}\log\left(\frac{2}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$. By Lemma E.5, we have that
${\mathbb{P}}\left(\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|>\epsilon\right)=$ ${\mathbb{P}}\left(|U^{2}-1|>n\epsilon\right)\leq 2\exp\left(-\frac{n^{2}\epsilon^{2}}{8}\right)$
where $U^{2}\sim\chi^{2}_{1}$ is a chi-square random variable. Hence, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|\leq\sqrt{\frac{8}{n^{2}}\log\left(\frac{2}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.11.
Let $z_{1},\cdots,z_{n},u_{1},\cdots,u_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|<\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$.
$\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)=\frac{1}{n}\left(\frac{\sum_{i=1}^{n}z_{i}}{\sqrt{n}}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{\sqrt{n}}\right)=\frac{1}{n}ZU,$
where $Z,U\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Observe that
$ZU=\frac{1}{4}(Z+U)^{2}-\frac{1}{4}(Z-U)^{2}\sim\frac{1}{2}Q-\frac{1}{2}R,$
where $Q:=\frac{1}{2}(Z+U)^{2},R:=\frac{1}{2}(Z-U)^{2}\sim\chi_{1}^{2}$ are i.i.d Chi-square distributed with 1 degree of freedom (since $\frac{1}{\sqrt{2}}(Z+U)$ and $\frac{1}{\sqrt{2}}(Z-U)$ are i.i.d standard Gaussian). Hence, by Lemma E.5 and a union bound, with probability at least $1-\delta$,
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|=$ $\left|\frac{1}{2}Q-\frac{1}{2}R\right|\leq\frac{1}{2}\left|Q-1\right|+\frac{1}{2}\left|R-1\right|$
$\leq$ $\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.12.
Let $A,B$ denote two PSD matrices such that $A\succeq B$. Then, their inverse $B^{-1}\succeq A^{-1}$.
Proof.
It suffices to show that $\sqrt{A}B^{-1}\sqrt{A}\succeq I$. Since $\sqrt{A}B^{-1}\sqrt{A}=(\sqrt{A}B^{-\frac{1}{2}})(\sqrt{A}B^{-\frac{1}{2}})^{\top}$ shares the same eigenvalues with $(\sqrt{A}B^{-\frac{1}{2}})^{\top}(\sqrt{A}B^{-\frac{1}{2}})=B^{-\frac{1}{2}}AB^{-\frac{1}{2}}\succeq I$, which concludes the proof. ∎

### v3

@@
-An Optimism-based Approach to Online Evaluation of Generative Models
+A Multi-Armed Bandit Approach to Online Selection and Evaluation of Generative Models
 Abstract
-Existing frameworks for evaluating and comparing generative models typically target an offline setting, where the evaluator has access to full batches of data produced by the models.
-However, in many practical scenarios, the goal is to identify the best model using the fewest generated samples to minimize the costs of querying data from the models.
-Such an online comparison is challenging with current offline assessment methods.
-In this work, we propose an online evaluation framework to find the generative model that maximizes a standard assessment score among a group of available models.
-Our method uses an optimism-based multi-armed bandit framework to identify the model producing data with the highest evaluation score, quantifying the quality and diversity of generated data.
-Specifically, we study the online assessment of generative models based on the Fréchet Inception Distance (FID) and Inception Score (IS) metrics and propose the FID-UCB and IS-UCB algorithms leveraging the upper confidence bound approach in online learning.
-We prove sub-linear regret bounds for these algorithms and present numerical results on standard image datasets, demonstrating their effectiveness in identifying the score-maximizing generative model.
+Existing frameworks for evaluating and comparing generative models consider an offline setting, where the evaluator has access to large batches of data produced by the models.
+However, in practical scenarios, the goal is often to identify and select the best model using the fewest possible generated samples to minimize the costs of querying data from the sub-optimal models.
+In this work, we propose an online evaluation and selection framework to find the generative model that maximizes a standard assessment score among a group of available models.
+We view the task as a multi-armed bandit (MAB) and propose upper confidence bound (UCB) bandit algorithms to identify the model producing data with the best evaluation score that quantifies the quality and diversity of generated data.
+Specifically, we develop the MAB-based selection of generative models considering the Fréchet Distance (FD) and Inception Score (IS) metrics, resulting in the FD-UCB and IS-UCB algorithms.
+We prove regret bounds for these algorithms and present numerical results on standard image datasets.
+Our empirical results suggest the efficacy of MAB approaches for the sample-efficient evaluation and selection of deep generative models.
+The project code is available at https://github.com/yannxiaoyanhu/dgm-online-eval.
 1 INTRODUCTION
 Deep generative models have achieved astonishing results across a wide array of machine learning datasets.
-Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) [17] and Inception Score (IS) [37].
+Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) (Heusel et al.,, 2017; Stein et al.,, 2023) and Inception Score (IS) (Salimans et al.,, 2016).
 Due to the growing applications of generative models to various learning tasks, the machine learning community has continuously adapted evaluation methodologies to better suit the characteristics of the newly introduced applications.
@@
 In particular, generating a large batch of high-resolution image or video data could be expensive and hinder the application of existing evaluation scores for ranking generative models.
-In this work, we focus on the online evaluation of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data.
+In this work, we focus on the online evaluation and selection of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data.
 By limiting the number of generated samples in the assessment process, online evaluation can save on the costs associated with producing large batches of samples.
 Additionally, online evaluation can significantly reduce the time and computational expenses required to identify a well-performing model in assessing a large group of generative models.
-To measure the performance of an online evaluation algorithm, we use the regret notion widely-used in the online learning literature.
+Figure 1 shows an example of FD-based evaluation of three pretrained models where the embeddings are extracted by InceptionNet.V3.
+While the offline evaluation procedure requires a large batch of generated data from each model, our online evaluation approach queries generated data from models adaptively and limits sample generation from the sub-optimal models, thus collecting samples from the best model more frequently (Figure 5).
+Figure 1: FID-based evaluation and selection among CIFAR10 generative models: The standard offline evaluation requires a large batch of data from every model.
+In contrast, our proposed online approach leverages the UCB multi-armed bandit strategy to identify the best model with fewer generations from the suboptimal models.
+To measure the performance of an online evaluation algorithm, we use the regret notion widely used in the online learning literature (Bubeck and Cesa-Bianchi,, 2012).
 In the online learning task, the evaluation algorithm selects a generative model and observes a mini-batch of samples generated by that model in each round.
@@
 Therefore, the learner’s regret quantifies the cost of generating samples from sub-optimal models.
-Our goal is to develop online learning algorithms that result in lower regret values in the assessment of generated data.
-To address the described online evaluation of generative models, we utilize the multi-armed bandit framework.
-In a standard multi-armed bandit problem, the online learner seeks to identify the arm with the highest expected value of a random score.
-While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FID metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
-By deriving concentration bounds for the FID and IS scores, we propose the optimism-based FID-UCB and IS-UCB algorithms to address the online assessment of generative models.
-These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FID and IS scores.
-Furthermore, we analytically bound the regret of the IS-UCB and FID-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
-We discuss the results of several numerical applications of FID-UCB and IS-UCB to standard image datasets and generative modeling frameworks.
+We aim to develop online learning algorithms that result in lower regret values in the assessment of generated data.
+To address the described online evaluation of generative models, we utilize the multi-armed bandit (MAB) framework.
+In a standard MAB setup, the online learner seeks to identify the arm with the highest expected value of a random score.
+While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FD metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
+By deriving concentration bounds for the FD and IS scores, we propose the optimism-based FD-UCB and IS-UCB algorithms to address the online assessment of generative models.
+These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FD and IS scores.
+Furthermore, we analytically bound the regret of the IS-UCB and FD-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
+We discuss the results of several numerical applications of FD-UCB and IS-UCB to standard image datasets and generative modeling frameworks.
 We compare the performance of these algorithms with the Greedy algorithm, which selects the model with the highest estimated score, and a Naive-UCB baseline that applies the UCB algorithm with a data-independent upper confidence bound.
 Our numerical results show a significant improvement in our proposed algorithms compared to the Greedy and Naive-UCB baselines.
-Additionally, we analyze the performance of FID-UCB with different image data embeddings, including InceptionNet.V3 [40], CLIP [7], and DINOv2 [33].
-Our empirical results indicate satisfactory performance under these standard embeddings.
+Additionally, FD-UCB can attain satisfactory performance under the standard image data embeddings, including InceptionNet.V3 (Szegedy et al.,, 2016), DINOv2 (Oquab et al.,, 2024), and CLIP (Radford et al.,, 2021).
+Our work demonstrates the effectiveness of MAB-based approaches to the evaluation and selection of generative models.
 The following summarizes the main contributions of this work:
-  • Proposing an online evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
-  • Developing the FID-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
-  • Proving sub-linear regret bounds for the proposed FID-UCB and IS-UCB algorithms.
-  • Demonstrating satisfactory empirical performance of FID-UCB and IS-UCB in comparison to the greedy and naive-UCB baseline algorithms.
+  • Proposing a MAB-based evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
+  • Developing the FD-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
+  • Proving sub-linear regret bounds for the proposed FD-UCB and IS-UCB algorithms.
+  • Demonstrating satisfactory empirical performance of FD-UCB and IS-UCB in comparison to the Greedy and Naive-UCB baseline algorithms.
 2 RELATED WORK
@@
 The evaluation of generative models has been extensively studied in the literature.
-Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic [3], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [5], and diversity/quality-based metrics such as Precision/Recall [36, 27], density and coverage [32], VENDI [13], and RKE [20].
-In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score [2], the FLD score [21], the Rarity score [16], and the KEN score [45] measuring the novelty of the generated samples.
+Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic (Arjovsky et al.,, 2017), Fréchet Inception Distance (FID) (Heusel et al.,, 2017) and Kernel Inception Distance (KID) (Bińkowski et al.,, 2018), and diversity/quality-based metrics such as Precision/Recall (Sajjadi et al.,, 2018; Kynkäänniemi et al.,, 2019), density and coverage (Naeem et al.,, 2020), Vendi (Friedman and Dieng,, 2023), RKE (Jalali et al.,, 2023), and scalable FKEA-Vendi Ospanov et al., (2024).
+In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score (Alaa et al.,, 2022), the FLD score (Jiralerspong et al.,, 2023), the Rarity score (Han et al.,, 2023), and the KEN score (Zhang et al., 2024b, ; Zhang et al., 2024a, ) measuring the novelty of the generated samples.
 In this paper, we primarily focus on FID and Inception Score, which have been frequently used for evaluating generative models.
 Role of embeddings in the quantitative evaluation results.
-Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 [40].
-However, [32] shows that such pretrained-embeddings could exhibit unexpected behaviors.
-Recently, several large pretrained models have been proposed, including DINOv2 [33] and CLIP [7].
-[39] shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models.
-In addition, [28] demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments.
+Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 (Szegedy et al.,, 2016).
+However, Naeem et al., (2020) shows that such pretrained embeddings could exhibit unexpected behaviors.
+Recently, several large pretrained models have been proposed, including DINOv2 (Oquab et al.,, 2024) and CLIP (Cherti et al.,, 2023).
+Stein et al., (2023) shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models.
+In addition, Kynkäänniemi et al., (2023) demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments.
 In this paper, we provide the numerical results for the mentioned embeddings extracted by different pretrained models.
@@
 Online learning is a sequential decision-making framework where an agent aims to minimize a cumulative loss function revealed to her sequentially.
-One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of [29, 41].
+One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of Lai and Robbins, (1985); Thompson, (1933).
 At each step, the agent chooses among several arms, each associated with a reward distribution, and aims to maximize a pre-specified performance metric.
-The primary concern of this body of literature considers maximizing the expected return [1, 4, 6].
-Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion [38, 47] in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards [44].
+The primary concern of this body of literature considers maximizing the expected return (Agrawal,, 1995; Auer,, 2003; Bubeck and Cesa-Bianchi,, 2012).
+Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion (Sani et al.,, 2012; Zhu and Tan,, 2020) in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards (Weinberger and Yemini,, 2023).
 However, to the best of our knowledge, the evaluation of generative models has not been exclusively studied in an online learning context.
+In a concurrent work, Rezaei et al., (2025) aim to maximize kernel-based evaluation scores by finding a mixture of generative models.
+In this work, we primarily focus on identifying the best (single) model.
 Online training of generative models.
-Several related works focus on training generative adversarial networks (GANs) [14] using online learning frameworks.
-[15] proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach.
-[10] shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs).
-Also, the recent paper [34] studies the no-regret behaviors of large language model (LLM) agents.
+Several related works focus on training generative adversarial networks (GANs) (Goodfellow et al.,, 2014) using online learning frameworks.
+Grnarova et al., (2018) proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach.
+Daskalakis et al., (2018) shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs).
+Also, the recent paper Park et al., (2024) studies the no-regret behaviors of large language model (LLM) agents.
 This reference proposes an unsupervised training loss, whose minimization could automatically result in known no-regret learning algorithms.
@@
 $\textup{IS}:=\exp\left\{\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[\textup{KL}(p_{Y|X_{\textup{g}}}\|p_{Y_{\textup{g}}})]\right\},$ (1)
-where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet [40], and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution.
+where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet (Szegedy et al.,, 2016), and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution.
 Further, we have that $\textup{IS}=\exp[I(X_{\textup{g}};Y_{\textup{g}})]=\exp[H(Y_{\textup{g}})-H(Y_{\textup{g}}|X_{\textup{g}})]$, where $I(\cdot;\cdot)$ is the mutual information, and $H(\cdot)$ is the Shannon entropy.
 A higher IS implies that the images generated by the model have higher diversity, since $p_{Y_{\textup{g}}}$ would be more uniformly distributed to increase $H(Y_{\textup{g}})$, and possess higher fidelity, because $p_{Y|X_{\textup{g}}}$ is closer to a one-hot vector to enforce a smaller $H(Y_{\textup{g}}|X_{\textup{g}})$.
-3.2 Fréchet Inception Distance
-Fréchet Inception Distance (FID) is another standard metric for evaluating generative models.
-Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$, which is extracted by the InceptionNet.V3 embedding.
-FID is the Fréchet distance [12] between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, defined as
-$\textup{FID}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}]$ (2)
-assuming that $p_{f(X_{\textup{g}})}={\mathcal{N}}(\mu_{\textup{g}},\Sigma_{\textup{g}})$ and $p_{f(X_{\textup{r}})}={\mathcal{N}}(\mu_{\textup{r}},\Sigma_{\textup{r}})$ are Gaussian distributions.
+3.2 Fréchet Distance
+Fréchet Distance (FD) (Dowson and Landau,, 1982) is another standard metric for evaluating generative models.
+Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$.
+The FD between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, which is computed by
+$\textup{FD}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}].$ (2)
+The well-known FID metric (Heusel et al.,, 2017) is computed as the FD where the image data feature is extracted by InceptionNet.V3 (Szegedy et al.,, 2016).
+Recently, Kynkäänniemi et al., (2023); Stein et al., (2023) propose computing the FD distance using the CLIP (Radford et al.,, 2021) and DINOv2 (Oquab et al.,, 2024) embeddings, respectively, to boost the score’s ranking consistency with human evaluations.
 4 ONLINE EVALUATION OF GENERATIVE MODELS
@@
 For each generator $g\in[G]$, we denote by $p_{g}\in\Delta({\mathcal{X}})$ its generative distribution over the space ${\mathcal{X}}$, which can be texts or images.
-Given an evaluation metric, e.g., FID or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$.
+Given an evaluation metric, e.g., FD or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$.
 The evaluation proceeds in $T$ steps.
@@
 In addition, the generative distribution typically lies in a high-dimensional space, which makes it difficult to estimate the score from limited data.
-Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FID requires the covariance matrix).
+Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FD requires the covariance matrix).
 Hence, analyzing the concentration properties of the score estimation would be challenging.
-Protocol 1 Online Evaluation of Generative Models
+Protocol 1 Online Evaluation and Selection of Generative Models
 […]
@@
 6:  end for
-5 Online Evaluation based on Fréchet Inception Distance
-In this section, we consider the online evaluation of generative models by Fréchet Inception Distance (FID).
-Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FID is computed by
-$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)
+Use of regret metric.
+The choice of regret metric 3 follows the standard online learning literature for bandit problems (Bubeck and Cesa-Bianchi,, 2012).
+Particularly, if the evaluator can attain a sub-linear regret, then the overall selection converges to the best model with high probability.
+Additionally, in the online evaluation task, it is possible that the best model has a very similar performance to that of the second-best model.
+In such a case, the regret would take the difference between the models into account, and in a case where the top-$k$ models have similar scores, generating samples from them would weighted more equally.
+On the other hand, if the best two models have significantly different scores, the regret would sharply grow by picking the suboptimal models.
+5 FRÉCHET DISTANCE-BASED ONLINE EVALUATION AND SELECTION
+Avg. Regret
+Figure 2: Online FD-based evaluation and selection among standard generative models: The $x$-axis is the number of online steps.
+At each step, the algorithm samples a batch of five generated images from the chosen model.
+The image data embeddings are extracted by CLIP (Cherti et al.,, 2023).
+Results are averaged over 20 trials.
+[…]
+In this section, we consider the online evaluation of generative models by Fréchet Distance (FD).
+Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FD is computed by
+$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)
 where
-$\widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top},\qquad\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i})$ (5)
-are the empirical covariance matrix and the mean vector, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3.
-It has been shown that the empirical FID (4) is biased differently depending on the generator, which makes it difficult to compare different generators for a fixed sample size [9].
-Hence, the FID-based evaluation is typically performed for a large batch of generated samples (e.g., $10\sim 50$k) to reduce the effect of the bias, which can be sample-inefficient and costly for a large number of generators.
-To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FID score.
-To this end, we first derive an optimistic FID score in the following theorem.
+$\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i}),\ \widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}$
+are the mean vector and the empirical covariance matrix, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3.
+The FD-based evaluation is typically performed on a large batch of generated samples (10-50 thousands) to reduce the estimation variance, which can be sample-inefficient and costly.
+To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FD score.
+To this end, we first derive an optimistic FD score in the following theorem.
 We defer the theorem’s proof to Appendix A.1.
-Theorem 5.1 (Generator-dependent optimistic FID score).
-Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive definite. Then, with probability at least $1-\delta$, we have
-$\widehat{\textup{FID}}_{g}^{n}=\widetilde{\textup{FID}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FID}_{g}.$ (6)
-Here, the bonus is given by
-${\mathcal{B}}_{g}^{n}:=$ $C^{n}_{g}\cdot\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (7)
-$+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\frac{2{\mathcal{I}}_{g}}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)},$
-where $C_{g}^{n}:=2(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})$, ${\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$, and $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$.
-Algorithm 1 FID-UCB for Generative Model Evaluation
+Theorem 1 (Optimistic FD score).
+Assume for any generator $g$, the (random) embedding $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$ follows a multivariate Gaussian, and the covariance matrix $\Sigma_{\textup{r}}$ of the real data is positive definite.
+Then, with probability at least $1-\delta$, we have
+$\widehat{\textup{FD}}_{g}^{n}=\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FD}_{g}$ (5)
+for any $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of the covariance matrix $\Sigma_{g}$.
+In addition, the bonus is given by
+${\mathcal{B}}_{g}^{n}:=$ $2\Delta_{\mu_{g}}^{n}\cdot\left(\Delta_{\mu_{g}}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{\Sigma_{g}}^{n}}$ (6)
+$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
+where $\Delta_{\mu_{g}}^{n}$ and $\Delta_{\Sigma_{g}}^{n}$ are defined in (23) and (24), respectively.
+Avg. Regret
+Figure 3: Online IS-based evaluation and selection among standard generative models: The $x$-axis is the number of steps.
+At each step, the algorithm samples a batch of five generated images from the chosen model.
+Results are averaged over 20 trials.
 […]
-1:  Initialize: $\widehat{\textup{FID}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in[G]$.
-2:  for step $t=1,2,\cdots,T$ do
-3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FID}}_{g}$.
-4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\in{\mathbb{R}}^{D}\sim g_{t}\}_{j=1}^{b}$.
-5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
-6:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
-7:     Compute the optimistic FID score
-$\widehat{\textup{FID}}_{g_{t}}\leftarrow\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}.$
-8:  end for
-We remark that the bonus term (7) is: 1) generator-dependent through parameters including ${\mathcal{I}}_{g}$, $\sigma_{g}^{2}$, $\textup{Tr}[\Sigma_{g}]$, and $\|\Sigma_{g}\|_{2}$, 2) dimension-free in the sense that the dimension $d$ only appears in the logarithmic term, which facilitates a sample-efficient evaluation.
-In practice, these parameters can be estimated from the queried data.
-Based on Theorem 5.1, we propose FID-UCB in Algorithm 1 as an FID-based online evaluation algorithm.
-At the beginning, the estimated FID scores are initialized to be negative infinity (line 1), and hence each generator is explored for at least one time.
-The evaluation proceeds iterative, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FID and queries a batch of images from the model (lines $3\sim 4$).
-Then, the estimated FID of generator $g_{t}$ is updated (lines $5\sim 7$).
-Next, we show that FID-UCB attains sub-linear regret bound.
-We defer the theorem’s proof to Appendix A.2.
-Theorem 5.2 (Regret of FID-UCB).
-Under the same conditions in Theorem A.1, with probability at least $1-\delta$, the regret of the UCB-FID algorithm after $T$ steps is bounded by
-$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$ (8)
-where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
-6 Online Evaluation based on Inception Score
+Remark 1 (Model-dependent parameters).
+The bonus term (6) involves model-dependent parameters $\textup{Tr}[\Sigma_{g}]$ and $\|\Sigma_{g}\|_{2}$ of the covariance matrix.
+Note that the presence of the norms of $\Sigma_{g}$ in the concentration bound for FD is inevitable, because, without any assumptions on the norm of $\Sigma_{g}$, the concentration bound cannot have a finite value.
+To address this, two approaches can be considered: 1) when the norm of the data embeddings is bounded, e.g., the normalized CLIP embeddings where $\|X_{g}\|_{2}\leq 1$, we can substitute $\textup{Tr}[\Sigma_{g}]$ and $\|\Sigma_{g}\|_{2}$ in the above bound with the embedding upper-bound norm.
+In this case, we can extend our analysis beyond multivariate Gaussians (Theorem 5 in the Appendix).
+2) On the other hand, when using the InceptionV3 and DINOv2 embeddings that have generally unbounded output norm, we can use the already generated data to estimate the parameters in the upper-bound.
+This approach avoids any offline estimation of the parameters required to have an upper-bound and preserves the online format of the FD-UCB algorithm, and conditioned that the estimated norms converge fast to the underlying value, will provide a correct solution.
+Our numerical results indicate that with only 50 samples, the norm terms in the bound will be relatively close to their actual values.
+See Part 3 in Appendix C for our experimental validation.
+Algorithm 1 FD-UCB
+0:  set of models ${\mathcal{G}}$, image embedding $f$, step $T\in{\mathbb{N}}_{+}$, $\mu_{\textup{r}}\in{\mathbb{R}}^{d}$ and $\Sigma_{\textup{r}}\in{\mathbb{S}}^{d}$ of real data, failure probability $\delta\leftarrow\delta/T$, batch size $b\in{\mathbb{N}}_{+}$, number of burn-in samples $N\geq\max_{g\in{\mathcal{G}}}\{4{\bm{r}}(\Sigma_{g})+\log(3/\delta)\}$.
+1:  Initialize: $\widehat{\textup{FD}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in{\mathcal{G}}$.
+2:  for model $g\in{\mathcal{G}}$ do
+3:     Generate $M$ samples $\{x_{0}^{j}\sim p_{g}\}_{j=1}^{N}$.
+4:     Update ${\mathcal{H}}_{g}\leftarrow{\mathcal{H}}_{g}\cup\{f(x_{0}^{j})\}_{j=1}^{N}$ and visitation $n_{g}\leftarrow N$.
+5:  end for
+6:  for step $t=1,2,\cdots,T$ do
+7:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FD}}_{g}$.
+8:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\sim g_{t}\}_{j=1}^{b}$.
+9:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
+10:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
+11:     Compute the optimistic FD score
+$\widehat{\textup{FD}}_{g_{t}}\leftarrow$ $-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}+\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}$
+$+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right],$
+where ${\mathcal{B}}^{n}_{gt}$ is given by Equation (6).
+12:  end for
+Based on Theorem 1, we propose FD-UCB in Algorithm 1 as an FD-based online evaluation algorithm.
+At the beginning, FD-UCB samples $N$ generated data from each model, after which the algorithm can compute an optimist FD score (5) for each model (lines $2$-$5$).
+The evaluation proceeds iteratively, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FD and queries a batch of images from the model (lines $7$-$8$).
+Then, the estimated FD of generator $g_{t}$ is updated (line $11$).
+It can be shown that FD-UCB attains sub-linear regret bound, which is formalized in Theorem 3 in Appendix A.2.
+6 INCEPTION SCORE-BASED ONLINE EVALUATION AND SELECTION
 In this section, we focus on evaluating generative models by Inception score (IS), which is given by $\textup{IS}=\exp[I(X_{g};Y_{g})]$, where $X_{g}\sim p_{g}$ is the generated image, and $Y_{g}$ is the class assigned by the InceptionNet.V3.
-Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, the empirical IS is computed by
-$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\},$ (9)
-where
-$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),\ \widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (10)
-are the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$ ($d=1000$ for InceptionNet.V3).
+Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, we denote by
+$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),$ (7)
+$\widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (8)
+the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$, respectively. Then, the empirical IS is computed by
+$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\}.$ (9)
 For any $j\in[d]$, let $\widehat{V}^{n}(p_{Y|X_{g}}[j]):={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ denote the sample variance for the probability that the generated sample is assigned to the $j$-th class.
@@
 where $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ is a $d$-dimensional vector whose $j$-th element is given by
-$\bm{\epsilon}^{n}_{g}[j]:=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$ (12)
+$\bm{\epsilon}^{n}_{g}[j]:=$ $\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}$ (11)
+$+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
 and $\textup{Clip}_{e^{-1}}(p,\bm{\epsilon})$ is the following element-wise operator
 $\begin{cases}p[j]+\frac{e^{-1}-p[j]}{|e^{-1}-p[j]|}\bm{\epsilon}[j]&\textup{, if }|e^{-1}-p[j]|\geq\bm{\epsilon}[j]\\ e^{-1}&\textup{, otherwise.}\end{cases}$ (12)
-The optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n}):=-\sum_{j}\widehat{p}_{Y_{g}}^{n}[j]\cdot\log(\widehat{p}_{Y_{g}}^{n}[j])\geq H(Y_{g})$ with high probability (see Lemma B.1 in the Appendix B.3).
-Next, we derive a generator-dependent optimistic IS in the following theorem.
-We defer the theorem’s proof to Appendix B.1.
-Theorem 6.1 (Generator-dependent Optimistic IS).
+Particularly, the optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$ with high probability (see Lemma 11 in the Appendix), where we define
+${\mathcal{E}}(z):=-\sum_{j}z[j]\cdot\log(z[j])$
+for any $d$-dimensional vector $z\succ\bm{0}$. Next, we derive a generator-dependent optimistic IS in the following theorem. We defer the theorem’s proof to Appendix B.1.
+Theorem 2 (Optimistic IS).
 Let
-$\widehat{\textup{IS}}^{n}_{g}:=\exp\left\{{\mathcal{E}}\left(\widehat{p}_{Y_{g}}^{n}\right)-\widehat{H}^{n}(Y_{g}|X_{g})+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)\right\},$ (14)
-where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (10), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (11), parameter $C^{\prime}$ satisfies that $P(|H(Y_{g}|X_{g})|\leq C^{\prime})=1$, and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution.
+[…]
+$+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4d}{\delta}\right)}$
+$+\frac{7\log d}{3(n-1)}\log\left(\frac{4d}{\delta}\right)\Bigg{\}},$
+where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (8), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (10), and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution.
 Then, with probability at least $1-\delta$, we have that $\widehat{\textup{IS}}^{n}_{g}\geq\textup{IS}_{g}$.
+Figure 4: Online IS-based evaluation and selection among variance-controlled FFHQ models: IS-UCB can identify models that generate images with more diversity. Results are averaged over 20 trials.
+[…]
 […]
@@
 5:     Update
-$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},\ \widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
+$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},$
+$\widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
 6:     Update visitation $N_{g_{t}}\leftarrow N_{g_{t}}+b$.
@@
 9:  end for
-Algorithm 2 IS-UCB for Generative Model Evaluation
-Based on Theorem 6.1, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm.
-In the following theorem, we derive the regret bound of the IS-UCB algorithm, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret.
-We present the theorem’s proof in Appendix B.2.
-Theorem 6.2 (Regret of IS-UCB).
-With probability at least $1-\delta$, the regret of the IS-UCB algorithm after $T$ steps is bounded by
-$\textup{Regret}(T)\leq\widetilde{O}\left(e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}+dG\Delta^{-2}\log d\right),$ (15)
-where $C:=O(de^{-1}+C^{\prime})$ is the upper bound of $\widehat{\textup{IS}}_{g}^{n}$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
+Algorithm 2 IS-UCB
+Based on Theorem 2, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm.
+Due to limited space in the main text, we derive the regret bound of the IS-UCB algorithm in Appendix B.2, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret.
 7 EXPERIMENTAL RESULTS
-Baseline methods.
-For both FID-based and IS-based evaluation, we compare the performance of our proposed FID-UCB and IS-UCB with the following two baselines, naive-UCB and the greedy algorithm.
-Naive-UCB is a simplification of FID-UCB and IS-UCB which is based on data-independent dimension-based concentration bounds, which do not exploit the generated data in evaluating the confidence bound.
-We present the exact estimators in Appendix D.
-On the other hand, the greedy algorithm always picks the generator with the lowest empirical FID (4) or the highest empirical IS (9).
-For all these algorithms, we initialize the estimated FID to be $\infty$- ∞ and the estimated IS to be +$\infty$+ ∞, and hence each generator will be explored at least once.
-[…]
+In this section, we present numerical results for the FD-UCB and IS-UCB algorithms.
+Our extensive experiments across standard image datasets, generative modeling frameworks, and image data embeddings demonstrate the effectiveness of the proposed online learning-based approaches to model evaluation and selection.
+Experimental details and additional results can be found in Appendix C.
+Baselines.
+For both FD-based and IS-based evaluation, we compare the performance of our proposed FD-UCB and IS-UCB with three baselines: Naive-UCB, Greedy algorithm, and Random algorithm.
+Naive-UCB is a simplification of FD-UCB and IS-UCB which replaces the generator-dependent variables in the bonus function with data-independent and dimension-based terms, which does not exploit the generated data in evaluating the confidence bound.
+On the other hand, the Greedy algorithm always picks the generator with the lowest empirical FD (4) or the highest empirical IS (9), and the Random algorithm selects the model randomly.
+For a fair comparison, we set the burn-in samples of FD-UCB to be zero.
+Additionally, for all these algorithms, each generator will be explored once at the beginning.
 Experimental setup and performance metrics.
-We report three performance metrics at each step $t\in[T]$: 1) total regret, i.e., $\textup{Regret}(t)$, 2) average regret, i.e., $(1/t)\cdot\textup{Regret}(t)$, and 3) optimal pick ratio, i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FID score or the highest empirical IS score for 50k generated images.
+We report two performance metrics at each step $t\in[T]$: 1) average regret (Avg.
+Regret), i.e., $(1/t)\cdot\textup{Regret}(t)$, and 2) optimal pick ratio (OPR), i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FD score or the highest empirical IS score for 50k generated images.
 For all the experiments, we use a batch size of 5, and the total evaluation step is $T=1,000$.
@@
 Results are averaged over 20 trials.
-Due to the limited space, we postpone parts of the results to Appendix C.
-7.1 Empirical Results of Online FID-based Evaluation
-Performance on pretrained generators under different embeddings.
-The results are summarized in Figure 1, where each row and each column correspond to one dataset and one performance metric, respectively.
-The results show that FID-UCB (blue) outperforms naive-UCB (orange) and the greedy algorithm (green).
-We observe that naive-UCB fails to distinguish between different generators and makes uniform selections on the CIFAR10 and the ImageNet datasets, while FID-UCB recognizes the optimal one.
-This shows that the generator-dependent optimistic FID (6) can better exploit the properties of the generator, which is key to the sample-efficient online evaluation.
-On the FFHQ dataset, all the generators attain very similar FID scores (the largest gap is less than 0.7), which may explain why FID-UCB has a similar performance to naive-UCB.
-We reran these experiments using DINOv2-ViT-L/14-based [33] and CLIP-based [7] embeddings.
-(The results are summarized in Figures 5 and 6 in Appendix C, respective.) The numerical results show that FID-UCB can maintain satisfactory performance over different embeddings.
-CIFAR10 (G=5)
-ImageNet (G=5)
-[…]
-Figure 1: Performance of FID-UCB, naive-UCB, and the greedy algorithm on various datasets. The embeddings of the images are extracted by the pretrained InceptionNet.V3.
-Performance on variance-controlled generators.
-We summarize the results in Figure 2.
-FID-UCB (blue) significantly outperforms naive-UCB and the greedy algorithm.
-The results for DINOv2-based and CLIP-based FID are provided in Figures 5 and 6 in Appendix C.
-We ran the evaluation on AFHQ Dog dataset, and the results are presented in Figure 7 in the Appendix.
-Furthermore, we display the generated images to visualize the quality and diversity of the samples queried by different algorithms.
-The results suggest that FID-UCB consistently queries images that are diverse and of high quality.
-In contrast, naive-UCB kept querying a proportion of images from the collapsed generator (the one with the smallest truncation parameter), and the greedy algorithm even stuck to it.
-FFHQ (G=20)
-Figure 2: Performance of FID-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
-Step 500
-[…]
-7.2 Results of Online IS-Based Evaluation
+Datasets and generators.
+We evaluate the above algorithms on standard image datasets, including CIFAR10 (Krizhevsky et al.,, 2009), ImageNet (Deng et al.,, 2009), FFHQ (Kazemi and Sullivan,, 2014), and AFHQ (Choi et al.,, 2020).
+For the first three datasets, we consider evaluation and selection among standard generative models, such as BigGAN-Deep (Brock et al.,, 2019), iDDPM-DDIM (Nichol and Dhariwal,, 2021), and Efficient-vdVAE (Hazami et al.,, 2022).
+Details of the chosen models can be found in Appendix C.
+Additionally, we consider variance-controlled generators for the FFHQ and AFHQ datasets, where we apply the standard truncation technique (Kynkäänniemi et al.,, 2019) to the pretrained StyleGAN2-ADA model (Karras et al.,, 2020) [^1] The repository can be found at github.com/NVlabs/stylegan2-ada-pytorch and synthesize $G=20$ models.
+Each model generates images from truncated random noises with parameters vary from $0.01$ to $0.1$.
+A small (large) truncation parameter can lead to generated samples with low (high) diversity but high (low) quality.
+7.1 Results of Online FD-based Evaluation and Selection
+FID-based evaluation and selection.
+We first report results for the setup in Figure 1, where we select among three standard generative models on the CIFAR10 dataset, including LOGAN (Wu et al.,, 2020), RESFLOW (Chen et al.,, 2019), and iDDPM-DDiM (Nichol and Dhariwal,, 2021).
+Our proposed FD-UCB can quickly identify the best model with fewer queries to the suboptimal models (Figure 5).
+Figure 5: Online FD-based evaluation and selection among three CIFAR10 models, including LOGAN, RESFLOW, and iDDPM-DDiM (Figure 1).
+The image data embeddings are extracted by InceptionV3.Net.
+Results are averaged over 20 trials.
+Performance on different embeddings.
+The results for CLIP embeddings are summarized in Figure 2, where each column and each row correspond to one dataset and one performance metric, respectively.
+The results show that FD-UCB outperforms Naive-UCB and the Greedy algorithm.
+We observe that naive-UCB converges to the best model much slower than FD-UCB, which suggests that the generator-dependent and data-driven bonus function can better exploit the properties of the generator, which is key to the sample-efficient online evaluation.
+We also test the InceptionNet.V3 and DINOv2 embeddings, and the results are summarized in Figures 7 and 8 in the Appendix.
+FD-UCB can maintain satisfactory performance over different embeddings.
+Performance on variance-controlled generators. We summarize the results on the FFHQ and AFHQ-Dog datasets in Figures 910, and 11 in the Appendix. FD-UCB consistently outperformed the baselines.
+Performance on video and audio data.
+We consider the Fréchet Video Distance (Unterthiner et al.,, 2019, FVD) and Fréchet Audio Distance (Kilgour et al.,, 2019, FAD) metrics for video and audio generation, respectively.
+We synthesize four arms (generators) utilizing the MSR-VTT (Xu et al.,, 2016) and Magnatagatune (Law et al.,, 2009) datasets.
+When an arm is selected, a video/audio clip is sampled from the dataset and perturbed by Gaussian noises with an arm-specific probability.
+Results on the audio data are summarized in Figure 6.
+Additional results can be found in Figure 12 in the Appendix.
+Figure 6: Online FAD-based evaluation and selection among four synthetic audio data.
+The embeddings are extracted by VGGish (Hershey et al.,, 2017).
+Results for the OPR metric are reported and averaged over 20 trials.
+7.2 Results of Online IS-Based Evaluation and Selection
 Performance on the pretrained generators.
@@
 IS-UCB (blue) attains significantly better performance than naive-UCB (orange) and the greedy algorithm (green) on CIFAR10 and FFHQ datasets.
-CIFAR10 (G=5)
-Figure 3: Performance of IS-UCB, naive-UCB, and the greedy algorithm on various datasets.
-ImageNet (G=5) [Uncaptioned image]
-[…]
 Performance on variance-controlled generators.
 The results are summarized in Figure 4.
-Generated samples are plotted in Figure 8 in the Appendix.
-The results on the AFHQ Dog dataset are presented in Figure 9 in the Appendix.
-FFHQ (G=20)
-Figure 4: Performance of IS-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
+The results suggest that FD-UCB consistently queries images that are diverse and of high quality.
+In contrast, Naive-UCB and the Greedy algorithm kept querying a proportion of images from the collapsed generators with smaller truncation parameters.
+The AFHQ-Dog dataset results are in the Appendix, Figure 13.
 8 CONCLUSION
 In this work, we studied an online learning problem aiming to identify the generative model with the best evaluation score among a set of models.
-We proposed the optimism-based FID-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models.
+We proposed the multi-armed bandit-based FD-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models.
 An interesting future direction to our work is to explore the application of other online learning frameworks, such as Thompson sampling, to address the online evaluation problem.
-Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FID-UCB and IS-UCB to text-based and video-based generative models will be interesting for future studies.
-Such an application only requires applying text and video-based embeddings, as in Video IS [35], FVD [42], and FTD [19].
-Appendix A Proofs in Section 5: FID Evaluation
-A.1 Proof of Theorem 5.1: Optimistic FID Score
+Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FD-UCB and IS-UCB to real text-based and video-based generative models will be interesting for future studies.
+Such an application only requires applying text and video-based embeddings, as in Video IS (Saito et al.,, 2020), FVD (Unterthiner et al.,, 2019), and FBD (Xiang et al.,, 2021).
+Also, the extension of our MAB framework to a contextual bandit algorithm for prompt-guided generative models, as studied by Hu et al., (2024), will remain for future studies.
+Acknowledgments
+The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No.
+4055164 and 4937054.
+The authors would also like to thank the anonymous reviewers for their constructive feedback and suggestions.
+Checklist
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. [Yes]
+    2. (b)
+      Complete proofs of all theoretical results. [Yes]
+    3. (c)
+      Clear explanations of any assumptions. [Yes]
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
+    4. (d)
+      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. [Yes]
+    2. (b)
+      The license information of the assets, if applicable. [Yes]
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. [Yes]
+    4. (d)
+      Information about consent from data providers/curators. [Yes]
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. [Not Applicable]
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
+Appendix A PROOFS IN SECTION LABEL:SEC:5: FD-BASED EVALUATION
+A.1 Proof of Theorem 1: Optimistic FD Score
 Proof.
-The proof of Theorem 5.1 is based on Theorem A.1 and Lemma A.2 in the Appendix.
-Specifically, Inequality (16) in Theorem A.1 in Appendix A.3 decomposes the estimation error of the empirical FID score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$).
-Next, Lemma A.2 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix.
-Combining ∎
-A.2 Proof of Theorem 5.2: Regret of FID-UCB
+The proof of Theorem 1 is based on Theorem 4 and Lemma 1 in the Appendix.
+Specifically, Inequality (18) in Theorem 4 in Appendix A.3 decomposes the estimation error of the empirical FD score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$): With probability at least $1-\frac{\delta}{3}$, it holds that
+$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|\leq$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$
+$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$
+Next, Lemma 1 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix: With probability at least $1-\frac{2\delta}{3}$, it holds that
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right)}=:\Delta_{\mu_{g}}^{n}$
+$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{\mu_{g}}^{n})^{2}=:\Delta_{\Sigma_{g}}^{n}$
+Therefore, let
+${\mathcal{B}}_{g}^{n}=2\Delta_{\mu_{g}}^{n}\cdot\left(\Delta_{\mu_{g}}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{\Sigma_{g}}^{n}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
+and we have that $\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{B}}_{g}^{n}\leq\textup{FD}_{g}$ holds with probability at least $1-\delta$, which concludes the proof. ∎
+A.2 Regret of FD-UCB
+Theorem 3 (Regret of FD-UCB).
+Under the same conditions in Theorem 1, with probability at least $1-\delta$, the regret of the FD-UCB algorithm with a batch size $b\in{\mathbb{N}}_{+}$ is bounded by
+$\widetilde{O}\left(\frac{D_{1}^{\mathcal{G}}\cdot G^{1/4}}{b^{1/4}}T^{3/4}+\frac{D_{2}^{\mathcal{G}}\cdot G^{1/2}}{b^{1/2}}T^{1/2}+\frac{D_{3}^{\mathcal{G}}\cdot G}{b}\log T+N\cdot G\right),$ (14)
+after running on $G:=|{\mathcal{G}}|$ models for $T\in{\mathbb{N}}_{+}$ steps, where $D_{1}^{\mathcal{G}}$ and $D_{2}^{\mathcal{G}}$ are model-dependent parameters defined in Equations (15)-(17), and the logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
 Proof.
 Recall that $g_{t}$ denotes the generator picked at the $t$-th step.
-For convenience, we denote by $\widehat{\textup{FID}}_{g_{t}}$ the optimistic FID of generator $g_{t}$ computed at step $t$.
-We denote by $n_{t}$ the number of images generated by model $g_{t}$ after $t$ steps.
-First, by Theorem A.2 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FID}_{g}\geq\widehat{\textup{FID}}_{g}$ for any step $t\in[T]$ and $g\in[G]$.
+For convenience, we denote by $\widehat{\textup{FD}}_{g_{t}}$ the optimistic FD of generator $g_{t}$ computed at step $t$.
+We denote by $n_{t}$ the number of images generated by model $g_{t}$ at the beginning of the $t$-th step.
+First, by Theorem 1 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FD}_{g_{t}}\geq\widehat{\textup{FD}}_{g_{t}}$ for any step $t\in[T]$ and $g\in[G]$.
 Hence, we have that
-$\textup{Regret}(T)=\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\textup{FID}^{*}\right)\leq\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widehat{\textup{FID}}_{g_{t}}\right).$
-By the definition of $\widehat{\textup{FID}}_{g_{t}}$, we further derive that
-$\textup{Regret}(T)\leq$ $\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widetilde{\textup{FID}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)\leq 2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}},$
-where ${\mathcal{B}}^{n}_{g}$ is the bonus function given by Equation (7). Therefore, the regret is further bounded by
-$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])\cdot T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$
-where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Here, we use the fact that
-$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{(2b)^{\alpha}}{1-\alpha}\sum_{g=1}^{G}(n_{g}(T))^{1-\alpha}\leq\frac{(2b)^{\alpha}}{1-\alpha}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
-$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq 2b\sum_{g=1}^{G}\log(n_{g}(T))\leq 2bG\log(T),$
-where $b$ is the batch size, and $n_{g}(T)$ is the number of images generated by model $g$ at the last step $T$. ∎
-A.3 Concentration of Empirical FID (4)
-Theorem A.1 (Concentration of empirical FID (4)).
-Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite, and we denote by $\Sigma_{\textup{r}}=Q_{\textup{r}}^{\top}\Lambda_{\textup{r}}Q_{\textup{r}}$ its eigendecomposition.
-Then, with probability at least $1-\frac{\delta}{3}$, we have that
-$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$ (16)
-$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}.$
+$\textup{Regret}(T)=O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\textup{FD}^{*}\right)\leq O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\widehat{\textup{FD}}_{g_{t}}\right),$
+where $\textup{FD}^{*}=\min_{g}\textup{FD}_{g}$, and the term $O(N\cdot G)$ corresponds to the regret incurred in the burn-in sampling phase.
+Further, by the definition of $\widehat{\textup{FD}}_{g_{t}}$, we further derive that
+$\textup{Regret}(T)\leq$ $O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\widetilde{\textup{FD}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)$
+$\leq$ $O(N\cdot G)+2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}}$
+$\leq$ $\widetilde{O}\left(N\cdot G+\sum_{t=1}^{T}\left(\Delta_{\mu_{g_{t}}}^{n_{t}}\cdot\left(\Delta_{\mu_{g_{t}}}^{n_{t}}+\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{\Delta_{\Sigma_{g_{t}}}^{n_{t}}}+\textup{Tr}[\Sigma_{g_{t}}]\sqrt{\frac{1}{n_{t}}}+\frac{\|\Sigma_{g_{t}}\|_{2}}{n_{t}}\right)\right),$
+where we utilize $\|\widehat{\mu}_{g_{t}}^{n_{t}}-\mu_{\textup{r}}\|_{2}\leq\Delta_{\mu_{g_{t}}}^{n_{t}}+\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}$ in the last inequality. Note that
+$\Delta_{\mu_{g_{t}}}^{n_{t}}=\sqrt{\frac{1}{n_{t}}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g_{t}}^{2}]\log\left(\frac{6T}{\delta}\right)}+8\|\Sigma_{g_{t}}\|_{2}\log\left(\frac{6T}{\delta}\right)\right)}=\widetilde{O}\left(n_{t}^{-1/2}\sqrt{\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}}\right)$
+and
+$\Delta_{\Sigma_{g_{t}}}^{n_{t}}=$ $20\kappa^{2}\|\Sigma_{g_{t}}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g_{t}})+\log(3T/\delta)}{n_{t}}}+(\Delta_{\mu_{g_{t}}}^{n_{t}})^{2}$
+$=$ $\widetilde{O}\left(n_{t}^{-1/2}\sqrt{\|\Sigma_{g_{t}}\|_{2}^{2}(\bm{r}(\Sigma_{g_{t}})+1)}+n_{t}^{-1}\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)\right).$
+Hence, the regret is further bounded by
+$\widetilde{O}\Bigg{(}N\cdot G+$ $\sum_{t=1}^{T}n_{t}^{-1/4}\cdot\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\left(\|\Sigma_{g_{t}}\|_{2}^{2}(\bm{r}(\Sigma_{g_{t}})+1)\right)^{1/4}$
+$+n_{t}^{-1/2}\cdot\left((\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}])\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)^{1/2}+\textup{Tr}[\Sigma_{g_{t}}]\right)$
+$+n_{t}^{-1}\cdot\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)\Bigg{)}.$
+Let
+$D_{1}^{\mathcal{G}}:=\max_{g}\left\{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\left(\|\Sigma_{g}\|_{2}^{2}(\bm{r}(\Sigma_{g})+1)\right)^{1/4}\right\},$ (15)
+$D_{2}^{\mathcal{G}}:=\max_{g}\left\{(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}])\left(\sqrt{\textup{Tr}[\Sigma_{g}^{2}]}+\|\Sigma_{g}\|_{2}\right)^{1/2}+\textup{Tr}[\Sigma_{g}]\right\},$ (16)
+$D_{3}^{\mathcal{G}}:=\max_{g}\left\{\sqrt{\textup{Tr}[\Sigma_{g}^{2}]}+\|\Sigma_{g}\|_{2}\right\}.$ (17)
+Then, we have that
+$\textup{Regret}(T)\leq\widetilde{O}\left(\frac{D_{1}^{\mathcal{G}}\cdot G^{1/4}}{b^{1/4}}T^{3/4}+\frac{D_{2}^{\mathcal{G}}\cdot G^{1/2}}{b^{1/2}}T^{1/2}+\frac{D_{3}^{\mathcal{G}}\cdot G}{b}\log T+N\cdot G\right)$
+where we use the fact that
+$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{2^{\alpha}}{b^{\alpha}(1-\alpha)}\sum_{g=1}^{G}(N_{g}(T))^{1-\alpha}\leq\frac{2^{\alpha}}{b^{\alpha}(1-\alpha)}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
+$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq\frac{2}{b}\sum_{g=1}^{G}\log(N_{g}(T))\leq\frac{2G}{b}\log T,$
+where $b$ is the batch size, and $N_{g}(T)$ is the number of picks of model $g$ at the last step $T$. Therefore, we conclude the proof. ∎
+A.3 Concentration of Empirical FD (4)
+The following theorem expresses the estimation error of the empirical FD by the concentration of the sample mean and covariance matrix, which facilitates the derivation of the bonus function.
+Theorem 4 (Concentration of empirical FD (4)).
+Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite. Then, with probability at least $1-\frac{\delta}{3}$, we have that
+$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|\leq$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$ (18)
+$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$
 Proof.
-Recall that for any generator $g$, we have that
-$\textup{FID}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
-$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
-(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  Note that
-$\left|\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}-\|\mu_{g}-\mu_{\textup{r}}\|_{2}\right|\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
-We derive
+For any generator $g$ and $n\in{\mathbb{N}}_{+}$, we have that
+$\textup{FD}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
+$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
+(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  We derive
 $\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}\right|$ (19)
-$\leq$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
-$\leq$ $(2\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
+$=$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
+$\leq$ $(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\text{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
 (2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.
-By Theorem A.3 in Appendix A.5, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
+By Lemma 2, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
 $\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}\left[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (20)
 (3) Bound $\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]$.  Note that
-$\left|\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]\right|$
-$=$ $\left|\mathbb{E}[\textup{Tr}[(f(X_{g})-\mu_{g})(f(X_{g})-\mu_{g})^{\top}]]-\frac{1}{n}\sum_{i=1}^{n}\textup{Tr}[(f(x^{i})-\widehat{\mu}_{g}^{n})((f(x^{i})-\widehat{\mu}_{g}^{n}))^{\top}]\right|$
-$=$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2}\right|$
-$\leq$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|.$ (19)
-Next, we bound the two terms in the last inequality. Let $Z_{g}:=f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. For the first term, we have that
-$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\|Z_{g}\|_{2}^{2}=(\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Sigma_{g}(\Sigma_{g}^{-\frac{1}{2}}Z_{g})=(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Lambda_{g}(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}),$
-where we denote by $\Sigma_{g}:=Q_{g}^{\top}\Lambda_{g}Q_{g}$ the eigendecomposition of $\Sigma_{g}$.
-Note that the random vector $Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}$ has distribution ${\mathcal{N}}(0,I)$.
-We have that
-$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\sum_{j=1}^{d}\lambda_{j}U_{j}^{2},$
-where $\lambda_{j}$’s are the eigenvalues of $\Sigma_{g}$, $U_{i}\sim{\mathcal{N}}(0,1)$ are independent standard normal, and hence $U_{i}^{2}$’s are independent Chi-square random variables.
-Therefore, we have that
-$\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|$
-$=$ $\left|\sum_{j=1}^{d}\lambda_{j}\left(\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-\mathbb{E}[U_{j}^{2}]\right)\right|$
-$\leq$ $\left(\sum_{j=1}^{d}\lambda_{j}\right)\cdot\max_{j=1,\cdots,d}\left|\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-1\right|$
-$\leq$ $\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)},$ (20)
-where the last inequality holds with probability at least $1-\frac{\delta}{3}$ by Lemma E.5. In addition, for the second term, we derive that
-$\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|$
-$=$ $\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})(\|f(x^{i})-\mu_{g}\|_{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right|$
-$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right)$
-$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right).$ (21)
+$\textup{Tr}[\widehat{\Sigma}_{g}^{n}]=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}\right]$ (21)
+$=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right]$
+$=$ $\textup{Tr}\left[\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}-(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right]$
+$=$ $\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}$
+Hence, we obtain
+$\left|\textup{Tr}[\Sigma_{g}]-\textup{Tr}[\widehat{\Sigma}_{g}^{n}]\right|\leq\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}$ (22)
+Note that $f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. By Lemma 4, with probability at least $1-\frac{\delta}{3}$, it holds that
+$\left|\frac{1}{n}\sum_{i=1}^{n}\|f(X^{i})-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]\right|\leq$ $\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$
 Putting all together.  Therefore, combining Inequalities (19), (20), and (22), with probability at least $1-\frac{\delta}{3}$, it holds that
-$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
-$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}},$
+$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|$
+$\leq$ $\underbrace{(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
+$+\underbrace{\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right)+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}$
+$=$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$
+$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
 which concludes the proof. ∎
 A.4 Concentration of Mean Vector and Covariance Matrix
-Lemma A.2 (Concentration of mean vector and covariance matrix).
-Under the same conditions in Theorem A.1, with probability at least $1-\frac{2\delta}{3}$, we have that
-$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}},\text{ where }{\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$ (22)
+Lemma 1 ($L$2-norm error for mean and covariance matrix).
+Under the same conditions in Theorem 4, with probability at least $1-\frac{2\delta}{3}$, we have that
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right)}=:\Delta_{\mu_{g}}^{n}$ (23)
 and
-$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2},$ (23)
-where $\sigma_{g}^{2}:=\max_{i\in[d]}\Sigma_{g}[i,i]$ is the largest variance for the elements in $f(X_{g})$, and $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
+$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{\mu_{g}}^{n})^{2}=:\Delta_{\Sigma_{g}}^{n}$ (24)
+for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
 Proof.
-1.
-Concentration of mean vector.
-Note that $\Sigma_{g}^{\frac{1}{2}}Z$ where $Z\sim{\mathcal{N}}(0,I)$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution, which is the same as $f(X_{g})-\mu_{g}$.
-We have that
-$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}(\Sigma_{g}^{\frac{1}{2}}Z_{i})\right\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}Z_{i}\right\|_{\Sigma_{g}},$
-where $\|y\|_{\Sigma_{g}}:=\sqrt{\langle y,\Sigma_{g}y\rangle}$ is the matrix norm. Let $\widehat{Z}:=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ denote the mean vector. We further obtain that
-$\left\|\widehat{Z}\right\|_{\Sigma_{g}}^{2}=\textup{Tr}\left((\widehat{Z})^{\top}\Sigma_{g}\widehat{Z}\right)=\textup{Tr}\left(\Sigma_{g}\widehat{Z}(\widehat{Z})^{\top}\right)=\textup{Tr}(\Sigma_{g}D)+\textup{Tr}(\Sigma_{g}M),$ (24)
-where $D:=\textup{diag}(\widehat{Z}(\widehat{Z})^{\top})$ and $M=\widehat{Z}(\widehat{Z})^{\top}-D$.
-For the first term, note that each diagonal element has the same distribution $\frac{1}{n^{2}}(\sum_{i=1}^{n}z_{i})^{2}$, where $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian.
-Then, with probability at least $1-\frac{\delta}{6}$, we have that
-$\textup{Tr}(\Sigma_{g}D)\leq\|D\|_{2}\textup{Tr}(\Sigma_{g})\leq\textup{Tr}(\Sigma_{g})\left(\frac{1}{n}\sqrt{32\log\left(\frac{12d}{\delta}\right)}\right),$ (25)
-where the first inequality holds by that $\Sigma_{g}$ is PSD, and the second inequality holds by Lemma E.10. For the second term,
-$\textup{Tr}(\Sigma_{g}M)=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot M_{j,l}$
-$=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)$
-$\leq$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\cdot\max_{j,l}\left|\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)\right|$
-where $z_{j,1},\cdots,z_{j,n},z_{l,1},\cdots,z_{l,n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. By Lemma E.11, with probability at least $1-\frac{\delta}{6}$, it holds that
-$\textup{Tr}(\Sigma_{g}M)\leq\left(\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\right)\frac{1}{n}\sqrt{8\log\left(\frac{24d}{\delta}\right)}$ (26)
-Combining Inequalities (25) and (26), we conclude that with probability at least $1-\frac{\delta}{3}$, it holds that
-$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|^{2}_{2}\leq\left(\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|\right)\left(\frac{2}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)$ (27)
+1. Concentration of squared $L$2-norm error.  Note that
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}=\underbrace{\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{(*): concentration}}+\underbrace{\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{expected L2-norm error}},$
+where $\widehat{\mu}_{g}^{n}-\mu_{g}\sim{\mathcal{N}}(0,\frac{\Sigma_{g}}{n})$.
+By Lemma 4, we have $\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]=\frac{\textup{Tr}[\Sigma_{g}]}{n}$, and with probability at least $1-\frac{\delta}{3}$, it holds that
+$\left|\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]\right|\leq\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right).$
 2. Concentration of sample covariance matrix. We have that
@@
 $=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}$
-$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2},$ (28)
-where the last inequality holds by the triangle inequality of the L2 norm of the matrix.
-Note that the zero-mean random vector $f(X_{g})-\mu_{g}$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution (under the assumption of FID).
-For the first term, we evoke Lemma E.7.
-Then, with probability at least $1-\frac{\delta}{3}$, it holds that
-$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}.$ (29)
+$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}\left((f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}-2(f(x^{i})-\mu_{g})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}+(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right)\right\|_{2}$
+$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}.$ (25)
+For the first term, note that $f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. Then, by Lemma 6, with probability at least $1-\frac{\delta}{3}$, it holds that
+$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}$ (26)
+for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ and $\kappa$ is a constant defined therein.
 For the second term, we have that
-$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2},$
+$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}\leq(\Delta_{\mu_{g}}^{n})^{2},$
 which concludes the proof. ∎
-A.5 Dimension-free Concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
-Theorem A.3 (Dimension-free concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
-Under the same conditions in Theorem A.1, we have that
+A.5 Concentration Bound for Norm-Bounded Embeddings
+We extend the concentration bound for norm-bounded embeddings.
+Theorem 5.
+Assume for any generator $g$, the (random) embedding $f(X_{g})$ is $L2$-norm bounded, i.e., $\|f(X_{g})\|_{2}\leq C$ for some positive number $C>0$, and the covariance matrix $\Sigma_{\textup{r}}$ of the real data is positive definite.
+Then, with probability at least $1-\delta$, we have
+$\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{C}}^{n}_{g}\leq\textup{FD}_{g},$
+where
+${\mathcal{C}}^{n}_{g}:=2(\Delta_{1}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\Delta_{1}^{n}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{2}^{n}}+4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{3}{\delta}\right)},$
+and $\Delta_{1}^{n}$ and $\Delta_{2}^{n}$ are defined in (28) and (29), respectively.
+Proof.
+The proof is similar to Theorem 1. Recall that for any generator $g$, we have that
+$\textup{FD}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
+$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
+(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  We derive
+$\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}\right|\leq(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+2\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
+(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.
+By Lemma 2, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
+$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}\left[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$
+(3) Bound $\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]$.  Note that
+$\left|\textup{Tr}[\Sigma_{g}]-\textup{Tr}[\widehat{\Sigma}_{g}^{n}]\right|\leq\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2}.$
+As $\|f(X_{g})\|_{2}\leq C$, we have $\|f(X_{g})-\mu_{g}\|_{2}^{2}\leq(\|f(X_{g})\|_{2}+\|\mu_{g}\|_{2})^{2}\leq(2C)^{2}$.
+By Hoeffding’s inequality, we have that with probability at least $1-\frac{\delta}{3}$, it holds that
+$\left|\frac{1}{n}\sum_{i=1}^{n}\|f(X^{i})-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]\right|\leq 4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}$
+Therefore, we have
+$\left|\textup{FD}_{g}-\widetilde{\textup{FD}}_{g}^{n}\right|\leq 2(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}+4C^{2}\sqrt{\frac{1}{2n}\log\left(\frac{3}{\delta}\right)}.$ (27)
+Next, we analyze the concentration error of the $L2$-norm error. Since
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}=\underbrace{\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{ concentration}}+\underbrace{\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]}_{\text{expected $L$2-norm error}},$
+the expected $L$2-norm error is given by
+$\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]=\frac{\textup{Tr}[\Sigma_{g}]}{n}.$
+Additionally, we have that $|\widehat{\mu}_{g}^{n}[i]-\mu_{g}[i]|\leq 2C$ for the $i$-th entry where $i\in[d]$.
+Hence, by Hoeffding’s inequality and a union bound over all dimensions, with probability at least $1-\frac{\delta}{3}$, it holds that
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}-\mathbb{E}[\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}]\leq\frac{2dC^{2}}{n}\log\left(\frac{6d}{\delta}\right).$
+Hence, we have
+$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2dC^{2}}{n}\log\left(\frac{6d}{\delta}\right)+\frac{\textup{Tr}[\Sigma_{g}]}{n}}:=\Delta_{1}^{n}$ (28)
+Further, we have that with probability at least $1-\frac{\delta}{3}$, it holds that
+$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{1}^{n})^{2}:=\Delta_{2}^{n}$ (29)
+for $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, which concludes the proof.
+A.6 Auxiliary Definitions and Lemmas
+Lemma 2 (Error decomposition for $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
+Under the same conditions in Theorem 4, we have that
 $\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (30)
 Proof.
-Let $\eta=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$. We have that
+Let $\eta:=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$ and $\Sigma_{\textup{r}}:=BB^{\top}$ denote the Cholesky decomposition of $\Sigma_{\textup{r}}$, where $B$ is invertible by the assumption $\Sigma_{\textup{r}}\succ\bm{0}$.
+Note that matrix $A\Sigma_{\textup{r}}=ABB^{\top}$ has the same set of eigenvalues with $B^{\top}AB$.
+We have that
 $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
-$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
+$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$ (Lemma 7)
 $\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B}+\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
 $\leq$ $\textup{Tr}\left[\sqrt{2\eta B^{\top}B}\right]=\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\eta},$
-where the first inequality holds by Lemma E.10, the second inequality holds by Lemma E.11 and the fact that both matrices $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B\succeq\bm{0},B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by Lemma E.10 and $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$.
+where the second inequality holds by Lemma 8 and the fact that $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B,B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$.
 By the same analysis, we can also derive that
@@
 which concludes the proof. ∎
-Appendix B Proofs in Section 6: IS Evaluation
-B.1 Proof of Theorem 6.1: Generator-Dependent Optimistic IS
+Definition 1 (Sub-Gaussian random variable).
+A random variable $X$ with mean $\mu=\mathbb{E}[X]$ is sub-Gaussian with (positive) parameter $\sigma$, which is denoted by $X\in\textup{SG}(\sigma)$, if
+$\mathbb{E}[\exp(s(X-\mu))]\leq\exp\left(\frac{s^{2}\sigma^{2}}{2}\right)$
+for all $s\in{\mathbb{R}}$.
+Definition 2 (Sub-Gaussian random vector).
+A random vector $X\in{\mathbb{R}}^{d}$ is a sub-Gaussian random vector with (positive) parameter $\sigma$ if $y^{\top}X\in\textup{SG}(\sigma)$.
+Particularly, the multivariate Gaussian $X\sim{\mathcal{N}}(0,\Sigma)\in\textup{SG}(\sqrt{\|\Sigma\|_{2}})$.
+Definition 3 (Sub-exponential random variable).
+A random variable $X$ with mean $\mu=\mathbb{E}[X]$ is sub-exponential with (non-negative) parameters $(\nu,\alpha)$, which denoted by $X\in\textup{SE}(\nu,\alpha)$, if
+$\mathbb{E}[\exp(s(X-\mu))]\leq\exp\left(\frac{s^{2}\nu^{2}}{2}\right)$
+for all $|s|<\frac{1}{\alpha}$. Moreover, we have
+${\mathbb{P}}[X-\mathbb{E}[X]\geq t]\leq\begin{cases}\exp(-\frac{t^{2}}{2\nu^{2}})&\textup{if }0\leq t\leq\frac{\nu^{2}}{\alpha},\\ \exp(-\frac{t}{2\alpha})&\textup{for }t>\frac{\nu^{2}}{\alpha}.\end{cases}$
+Equivalently, with probability at least $1-\delta$, it holds that
+$|X-\mathbb{E}[X]|\leq\max\left\{\sqrt{2\nu^{2}\log\left(\frac{2}{\delta}\right)},2\alpha\log\left(\frac{2}{\delta}\right)\right\}.$
+Lemma 3 (Weighted sum of sub-exponential r.v.’s).
+Let $X_{1},\cdots,X_{n}$ be $n$ independent random variables, where $X_{i}\in\textup{SE}(\nu_{i},\alpha_{i})$ is sub-exponential with mean $\mathbb{E}[X_{i}]=\mu_{i}$ and parameters $(\nu_{i},\alpha_{i})$.
+Then, for any $W=(w_{1},\cdots,w_{n})^{\top}\in{\mathbb{R}}^{n}_{+}$, the (non-negative) weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$ is sub-exponential with parameters $(\sqrt{\sum_{i=1}^{n}w_{i}^{2}\nu_{i}^{2}},\max_{i}\{w_{i}\cdot\alpha_{i}\})$.
 Proof.
-By Lemma B.2, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 6.1, it suffices to show that
-$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}-\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)$
-with probability at least $1-\frac{\delta}{2}$, which holds by Theorem E.4. Therefore, we conclude the proof. ∎
-B.2 Proof of Theorem 6.2: Regret of IS-UCB
+By Definition 3, we have
+$\mathbb{E}[\exp(s\cdot w_{i}(X_{i}-\mu_{i}))]\leq\exp\left(\frac{s^{2}w_{i}^{2}\nu_{i}^{2}}{2}\right)$
+for any $i\in[n]$ and $|s|\cdot w_{i}<\frac{1}{\alpha_{i}}$ Further, since $X_{1},\cdots,X_{n}$ are independent, it holds that
+$\mathbb{E}[\exp(s\cdot\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{i}))]\leq\prod_{i=1}^{n}\exp\left(\frac{s^{2}w_{i}^{2}\nu_{i}^{2}}{2}\right)=\exp\left(\frac{s^{2}\left(\sqrt{\sum_{i=1}^{n}w_{i}^{2}\nu_{i}^{2}}\right)^{2}}{2}\right)$
+for any $|s|<\min_{i}\frac{1}{w_{i}\cdot\alpha_{i}}$, which concludes the proof. ∎
+Lemma 4 (Squared L2-norm of multivariate Gaussian).
+Let $X\sim{\mathcal{N}}(\bm{0},\Sigma)$ be any $d$-dimensional multivariate Gaussian with zero mean.
+Then, the r.v.
+$\|X\|_{2}^{2}\sim\sum_{i=1}^{d}\lambda_{i}\chi^{2}_{1}$ is a weighted sum of i.i.d.
+chi-squared random variables with one degree of freedom, where $\{\lambda_{i}\}_{i=1}^{d}$ are the eigenvalues of the covariance matrix $\Sigma$.
+Further, let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\bm{0},\Sigma)$ denote i.i.d.
+samples.
+Then, with probability at least $1-\delta$, it holds that
+$\left|\frac{1}{n}\sum_{i=1}^{n}\|X_{i}\|_{2}^{2}-\mathbb{E}[\|X\|_{2}^{2}]\right|\leq\max\left\{\sqrt{\frac{8\cdot\textup{Tr}[\Sigma^{2}]}{n}\log\left(\frac{2}{\delta}\right)},\frac{8\|\Sigma\|_{2}}{n}\log\left(\frac{2}{\delta}\right)\right\},$ (31)
+where $\mathbb{E}[\|X\|_{2}^{2}]:=\sum_{i=1}^{d}\lambda_{i}=\textup{Tr}[\Sigma]$.
 Proof.
-By Theorem 6.1 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g}\geq\textup{IS}_{g}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (14).
+Let $\Sigma=Q\Lambda Q^{\top}$ be the eigendecomposition of the covariance matrix, where we define $\Lambda:=\textup{diag}(\lambda_{1},\cdots,\lambda_{d})$.
+Hence, we have $X\sim QZ$, where $Z\sim{\mathcal{N}}(\bm{0},\Lambda)$ is isotropic multivariate Gaussian.
+Hence, we derive
+$\|X\|_{2}^{2}=(QZ)^{\top}QZ=Z^{\top}Q^{\top}QZ=Z^{\top}Z=\sum_{i=1}^{d}\lambda_{i}\chi_{1}^{2}.$
+Since $\chi_{1}^{2}\in\textup{SE}(2,4)$ is sub-exponential with parameters $(2,4)$, by Lemma 3 (with $w_{i}=\lambda_{i}$), we have $\|X\|_{2}^{2}\in\textup{SE}(2\sqrt{\textup{Tr}[\Sigma^{2}]},4\|\Sigma\|_{2})$ and has mean $\mathbb{E}[\|X\|_{2}^{2}]=\textup{Tr}[\Sigma]$.
+We conclude the proof by Lemma 5.
+∎
+Lemma 5 (Bernstein-type bound for sub-exponential r.v.).
+Let $X_{1},\cdots,X_{n}$ be $n$ i.i.d. random variables, where each $X_{i}\in\textup{SE}(\nu,\alpha)$. Then, with probability at least $1-\delta$, it holds that
+$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mathbb{E}[X]\right|\leq\max\left\{\sqrt{\frac{2\nu^{2}}{n}\log\left(\frac{2}{\delta}\right)},\frac{2\alpha}{n}\log\left(\frac{2}{\delta}\right)\right\}$
+Proof.
+By Lemma 3, we have $\frac{1}{n}\sum_{i=1}^{n}X_{i}\in\textup{SE}(\frac{\nu}{\sqrt{n}},\frac{\alpha}{n})$, which concludes the proof. ∎
+Lemma 6 (Dimension-free concentration of sample covariance matrix (Lounici,, 2014; Koltchinskii and Lounici,, 2017; Zhivotovskiy,, 2022)).
+Let $X_{1},\cdots,X_{n}$ are i.i.d sub-Gaussian random vectors with zero mean and covariance matrix $\Sigma$. Let
+$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (32)
+denote the effective rank. Then, with probability at least $1-\delta$, the sample covariance matrix satisfies that
+$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\kappa^{2}\|\Sigma\|_{2}\left(\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}\right)$ (33)
+for $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$.
+Here, the parameter $\kappa$ is a constant such that $\|y^{\top}X\|_{\psi_{2}}\leq\kappa\sqrt{y^{\top}\Sigma y}$, where $\|Z\|_{\psi_{2}}:=\inf\{t>0:\mathbb{E}[\exp(Z^{2}/t^{2})]\leq 2\}$ is the sub-Gaussian norm.
+Lemma 7.
+Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
+$\textup{Tr}(\sqrt{A+B})\geq\max\{\textup{Tr}(\sqrt{A}),\textup{Tr}(\sqrt{B})\}$ (34)
+Proof.
+Since $A+B\succeq A$ and $A+B\succeq B$, by Courant-Fischer min-max theorem, it holds that $\lambda_{i}(A+B)\geq\lambda_{i}(A)$ and $\lambda_{i}(A+B)\geq\lambda_{i}(B)$ for their $i$-th largest eigenvalues, which concludes the proof.
+∎
+Lemma 8.
+Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
+$\textup{Tr}[\sqrt{A+B}]\leq\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}].$ (35)
+Proof.
+First, we assume that both matrices $A$ and $B$ are positive strictly definite. Then,
+$\textup{Tr}(\sqrt{A+B})=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}(A+B)(A+B)^{-\frac{1}{4}}]$
+$=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}A(A+B)^{-\frac{1}{4}}]+\textup{Tr}[(A+B)^{-\frac{1}{4}}B(A+B)^{-\frac{1}{4}}]$
+$=$ $\textup{Tr}[\sqrt{A}(A+B)^{-\frac{1}{2}}\sqrt{A}]+\textup{Tr}[\sqrt{B}(A+B)^{-\frac{1}{2}}\sqrt{B}]$
+$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
+where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma 9). For general PSD matrices $A$ and $B$, note that
+$\textup{Tr}(\sqrt{A+B+\epsilon I})\leq\textup{Tr}(\sqrt{A+\epsilon I})+\textup{Tr}(\sqrt{B+\epsilon I})$
+holds for any $\epsilon>0$ by our previous analysis. Since the trace map is continuous, ∎
+Lemma 9.
+Let $A,B$ denote two PSD matrices such that $A\succeq B$. Then, their inverse $B^{-1}\succeq A^{-1}$.
+Proof.
+It suffices to show that $\sqrt{A}B^{-1}\sqrt{A}\succeq I$.
+Since $\sqrt{A}B^{-1}\sqrt{A}=(\sqrt{A}B^{-\frac{1}{2}})(\sqrt{A}B^{-\frac{1}{2}})^{\top}$ shares the same eigenvalues with $(\sqrt{A}B^{-\frac{1}{2}})^{\top}(\sqrt{A}B^{-\frac{1}{2}})=B^{-\frac{1}{2}}AB^{-\frac{1}{2}}\succeq I$, which concludes the proof.
+∎
+Appendix B PROOFS IN SECTION LABEL:SEC:6: IS-BASED EVALUATION
+B.1 Proof of Theorem 2: Generator-Dependent Optimistic IS
+Proof.
+By Lemma 10, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 2, it suffices to show that
+$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4d}{\delta}\right)}-\frac{7\log d}{3(n-1)}\log\left(\frac{4d}{\delta}\right)$
+with probability at least $1-\frac{\delta}{2}$, which holds by Theorem 12 and the fact that $H_{g}(Y_{g}|X_{g})\leq\log d$. Therefore, we conclude the proof. ∎
+B.2 Regret of IS-UCB
+Theorem 6 (Regret of IS-UCB).
+With probability at least $1-\delta$, the regret of the IS-UCB algorithm after $T$ steps is bounded by
+$\textup{Regret}(T)\leq\widetilde{O}\left(G\Delta^{-2}\log d+e^{C}\cdot\left((dm+\log d)\cdot\sqrt{T}+\log T\right)\right),$ (36)
+where $C=\widetilde{O}(d)$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
+Proof.
+By Theorem 2 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g_{t}}\geq\textup{IS}_{g_{t}}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (13).
 For convenience, we denote by $\widehat{\textup{IS}}_{g_{t}}$ the optimistic IS of generator $g_{t}$ computed at the $t$-th step.
@@
 $\textup{IS}^{*}-\textup{IS}_{g_{t}}\leq\widehat{\textup{IS}}_{g_{t}}-\textup{IS}_{g_{t}}=e^{{\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}}-\textup{IS}_{g_{t}},$
-where
-${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right).$
+where we denote
+${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n_{t}}(H(Y_{g}|X_{g}))}{n_{t}}\log\left(\frac{4Td}{\delta}\right)}+\frac{7\log d}{3(n_{t}-1)}\log\left(\frac{4Td}{\delta}\right)$
+for convenience.
 Let ${\mathcal{T}}_{1},{\mathcal{T}}_{2}\subset[T]$ be two disjoint set of steps such that ${\mathcal{T}}_{2}=[T]\backslash{\mathcal{T}}_{1}$.
@@
 $\leq$ $O\left(e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(m\cdot\sum_{j=1}^{d}\bm{\epsilon}_{g_{t}}[j]+{\mathcal{B}}_{g_{t}}\right)\right)$
-$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left(\max\left\{dm,\frac{d\log d}{2}\right\}\sqrt{\frac{1}{N_{g_{t}}}}+\frac{C^{\prime}}{N_{g_{t}}}\right)\right),$ (31)
-where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=O(de^{-1}+C^{\prime})$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$.
+$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left((dm+\log d)\sqrt{\frac{1}{N_{g_{t}}}}+\frac{\log d}{N_{g_{t}}}\right)\right),$ (37)
+where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=\widetilde{O}(d)$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$.
 For the second part, note that the number of steps where at least one element of $\widetilde{p}_{Y_{g_{t}}}$ is clipped to at most $\widetilde{O}(G\Delta^{-2})$, where $\Delta=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$.
 Therefore, we have that
-$\textup{Regret}(T)\leq\widetilde{O}\left(dG\Delta^{-2}\log d+e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}\right),$
+$\textup{Regret}(T)\leq\widetilde{O}\left(G\Delta^{-2}\log d+e^{C}\cdot\left((dm+\log d)\cdot\sqrt{T}+\log T\right)\right),$
 which concludes the proof. ∎
-B.3 Optimistic marginal class distribution
-Theorem B.1 (Optimistic marginal class distribution).
+B.3 Data-dependent optimistic marginal class distribution
+Lemma 10 (Optimistic marginal class distribution).
+Let $x^{1},\cdots,x^{n}\sim p_{g}$ be $n$ generated images from generator $g$. Define $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ denote the error vector whose $j$-th element is given by
+$\bm{\epsilon}^{n}_{g}[j]=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
+where $\widehat{V}^{n}(p_{Y|X_{g}}[j])={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ is the empirical variance for the $j$-th class density.
+Then, with probability at least $1-\frac{\delta}{2}$, we have that
+$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g}\right)$
+satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
+Proof.
+It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ for all $j\in[d]$ with probability at least $1-\frac{\delta}{2}$.
+We evoke Lemma 12 and conclude the proof.
+∎
+B.4 Auxiliary Definitions and Lemmas
+Lemma 11 (Optimistic marginal class distribution).
 Let $\widetilde{p}^{n}_{Y_{g}}:=\frac{1}{n}\sum_{i=1}^{n}p_{Y_{g}|x^{i}}$ denote the empirical marginal class distribution. Let
@@
 ∎
-B.4 Data-dependent optimistic marginal class distribution
-Lemma B.2 (Optimistic marginal class distribution).
-Let $x^{1},\cdots,x^{n}\sim p_{g}$ be $n$ generated images from generator $g$. Define $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ denote the error vector whose $j$-th element is given by
-$\bm{\epsilon}^{n}_{g}[j]=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
-where $\widehat{V}^{n}(p_{Y|X_{g}}[j])={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ is the empirical variance for the $j$-th class density.
-Then, with probability at least $1-\frac{\delta}{2}$, we have that
-$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g}\right)$
-satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
-Proof.
-It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ with probability at least $1-\frac{\delta}{2}$. We evoke Theorem E.4 and conclude the proof. ∎
-Appendix C Additional Experimental Results
-C.1 Results of FID-Based Online Evaluation
-1. Embeddings Extracted by DINOv2 [33].
-We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by DINOv2-ViT-L/14.
-The pretrained weights are downloaded from the official repository (licensed under Apache License 2.0).
-[^3] https://github.com/facebookresearch/dinov2/tree/main The results are summarized in Figure 5.
-CIFAR10 (G=5)
-ImageNet (G=5)
-FFHQ (G=5)
-[…]
-Figure 5: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators.
-The embeddings of the images are extracted by the pretrained DINOv2-ViT-L/14 model.
-2. Embeddings Extracted by CLIP [7].
-We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by CLIP.
-The pretrained weights are downloaded from the official repository.
-[^4] https://github.com/mlfoundations/open_clip.
-Copyright (c) 2012-2021 Gabriel Ilharco, Mitchell Wortsman, Nicholas Carlini, Rohan Taori, Achal Dave, Vaishaal Shankar, John Miller, Hongseok Namkoong, Hannaneh Hajishirzi, Ali Farhadi, Ludwig Schmidt.
-The results are summarized in Figure 6.
-We note that in the FFHQ dataset, FID-UCB attains a low optimal pick ratio as the optimal generator and the second-best one attains almost the same CLIP-based FID.
-CIFAR10 (G=5)
-ImageNet (G=5)
-FFHQ (G=5)
-[…]
-Figure 6: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators.
-The embeddings of images are extracted by the pretrained CLIP model.
-3. Results on AFHQ Dog dataset.
-The results are summarized in Figure 7.
-AFHQ Dog (G=20)
-Figure 7: Performance of FID-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ Dog dataset. The embeddings of images are extracted by InceptionNet.V3.
-C.2 Results of IS-Based Online Evaluation
-1. Diversity and quality of the generated samples.
-We examine the diversity and quality of the generated samples queried by IS-UCB, naive UCB, and the greedy algorithm (i.e., experiment in Figure 4). The results are summarized in Figure 8.
-Step 500
-[…]
-Figure 8: Generated images queried by IS-UCB, naive UCB, and the greedy algorithm.
-2. Results on AFHQ Dog dataset.
-The results are summarized in Figure 9.
-AFHQ Dog (G=20)
-Figure 9: Performance of IS-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ dataset.
-Appendix D Additional Experimental details
-1. Additional experiment setups.
-All the experiments are run on a single Nvidia GeForce RTX 3090 GPU.
-For FID-based evaluation, each trial takes at most 2 hours to finish, depending on the dataset and the embedding pretrained model.
-For IS-based evaluation, each trial takes less than 30 minutes to finish.
-2. List of pretrained generative models/generated image data.
-For the CIFAR10 dataset, we compare pretrained models including SNGAN, ACGAN, BigGAN, ContraGAN, and BigGAN-CR.
-For the ImageNet dataset, we compare pretrained models including StyleGAN3, SNGAN, BigGAN, ContraGAN, and ReACGAN.
-The pretrained weights are downloaded from the StudioGAN repository [22].
-For the FFHQ dataset, we compare StyleSwin, StyleNAT, StyleGAN2-ADA, StyleGAN-XL, and LDM, where we utilize the generated image datasets downloaded from the dgm-eval repository [39].
-3. Implementation details of naive UCB for FID-based evaluation.
-In FID-based evaluation, naive UCB upper bounds the error in estimating the mean vector by
-$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\sqrt{\sum_{j=1}^{d}(\widehat{\mu}_{g}^{n}[j]-\mu_{g}[j])^{2}}\leq\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{2d}{\delta}\right)}$
-which holds by the assumption of FID score that $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$, with probability at least $1-\delta$.
-In addition, the error in estimating the covariance matrix is bounded by
-$\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}+\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}\right)^{2}\right)$
-which follows from the exact same derivation in Inequality (28) and holds with probability at least $1-\delta$ (see Lemma E.6).
-Compared to the generator-dependent error bounds utilized by FID-UCB (i.e., Inequalities (27) and (29)), these bounds are dimension-dependent and do not fully exploit the properties of the generator.
-Finally, by the exact same analysis in the proof of Theorem 5.1, the following estimated FID
-$\widehat{\textup{FID}}_{g}^{n}=$ $\widetilde{\textup{FID}}_{g}^{n}-C_{g}^{n}\cdot\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{6d}{\delta}\right)}-\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (32)
-$-\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}\right)^{2}\right)+\frac{d\sigma_{g}^{2}}{n}\log\left(\frac{6d}{\delta}\right)\right)}$
-lower bounds $\textup{FID}_{g}$ with probability at least $1-\delta$.
-4. Implementation details of naive UCB for IS-based evaluation.
-In IS-based evaluation, naive UCB uses Hoeffding-type bonus terms.
-Specifically, the error vector in defining the optimistic marginal class distribution is given by $e_{g}^{n}$, whose $j$-th element is defined as
-$e_{g}^{n}[j]:=\sqrt{\frac{1}{2n}\log\left(\frac{4d}{\delta}\right)},$
-which ensures that ${\mathcal{E}}(\textup{Clip}_{e^{-1}}(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}))\geq H(Y_{g})$ with probability at least $1-\frac{\delta}{2}$.
-In addition, the concentration error of $\widehat{H}^{n}(Y_{g}|X_{g})$ is bounded by
-$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}$
-which holds with probability at least $1-\frac{\delta}{2}$. Therefore, the estimated IS is given by
-$\widehat{\textup{IS}}_{g}^{n}:=\exp\left\{{\mathcal{E}}\left(\textup{Clip}_{e^{-1}}\left(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}\right)\right)-\widehat{H}^{n}(Y_{g}|X_{g})+C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}\right\}$ (33)
-By the exact same analysis in the proof of Theorem 6.1, it can be shown that $\widehat{\textup{IS}}_{g}^{n}\geq\textup{IS}_{g}$ with probability at least $1-\delta$.
-5. Additional implementation details of FID-UCB
-To compute the bonus term (7), we need to estimate the following parameters: 1) $\|\mu_{g}-\mu_{\textup{r}}\|_{2}$: we set it to be a constant, e.g., 0.5; 2) ${\mathcal{I}}_{g}$: we estimate this term using the truncated covariance matrix with parameter $\tau$, which is denoted by $\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})$.
-The $(i,j)$-th element of the truncated covariance matrix is $\widehat{\Sigma}_{g}^{n}[i,j]$ if $\widehat{\Sigma}_{g}^{n}[i,j]\geq\tau$ and is zero otherwise.
-In the experiments, we observe that a truncation parameter that is $5\sim 10\%$ of the operator norm $\|\Sigma_{\textup{r}}\|_{2}$, i.e., $0.05\cdot\|\Sigma_{\textup{r}}\|_{2}\leq\tau\leq 0.1\cdot\|\Sigma_{\textup{r}}\|_{2}$, attains satisfying performance; 3) $\|\Sigma_{g}\|_{2}$: we estimate this term by $\max_{j\in[d]}(\widehat{\sigma}_{j}^{n})^{2}+O(\sqrt{\log(1/\delta)/n})$, where $(\widehat{\sigma}_{j}^{n})^{2}$ is the sample variance for the $j$-th entry of the embedding $f(X_{g})$.
-This avoids the computation for the eigendecomposition of the sample covariance matrix $\widehat{\Sigma}_{g}^{n}$.
-The intuition is that when $\Sigma_{g}$ is diagonal, the operator norm $\|\Sigma_{g}\|_{2}$ coincides with the largest per-entry variance; 4) $\textup{Tr}[\Sigma_{g}]$: while a direct approach is to estimate it by $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$, we found that a better performance can be attained by using $O(\textup{Tr}[\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})\cdot\Sigma_{\textup{r}}])$, where we utilize the fact that $\|\Sigma_{\textup{r}}^{-1}\|_{2}\cdot\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]\geq\textup{Tr}[\Sigma_{g}]\geq\frac{\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]}{\|\Sigma_{\textup{r}}\|_{2}}$.
-Appendix E Auxiliary Lemmas
-Lemma E.1 (Concentration of Gaussian random variable).
-Let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$. Then, it holds that
-${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|>\epsilon\right)<\exp\left(-\frac{n\epsilon^{2}}{2\sigma^{2}}\right)$
-Equivalently, with probability at least $1-\delta$, we have that
-$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|\leq\sqrt{\frac{\sigma^{2}}{2n}\log\left(\frac{2}{\delta}\right)}.$
-Lemma E.2 (Hoeffding’s inequality).
-Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d. random variables with values in $[0,1]$ and let $\delta>0$. Then with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
-$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{1}{2n}\log\left(\frac{1}{\delta}\right)}.$ (34)
-A drawback of Hoeffding’s inequality is that the confidence interval is distribution-agnostic, and always scales with order $O(\sqrt{1/n})$.
-For distribution with small variance, the following [18, Bennett’s inequality] provides a tighter concentration error.
-Lemma E.3 (Bennett’s inequality).
-Under the conditions of Theorem E.2, with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
-$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2{\mathbb{V}}Z}{n}\log\left(\frac{1}{\delta}\right)}+\frac{1}{3n}\log\left(\frac{1}{\delta}\right),$ (35)
-where ${\mathbb{V}}Z$ is the variance ${\mathbb{V}}=\mathbb{E}(Z-\mathbb{E}Z)^{2}$.
-When the exact variance is unknown, [31] provide a purely data-dependent bound with similar properties as Bennett’s inequality.
-Lemma E.4 (Restatement of [31, Theorem 4]).
-Under the conditions of Theorem E.2, we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
+Lemma 12 (Restatement of (Maurer and Pontil,, 2009, Theorem 4)).
+Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d.
+random variables with values in $[0,1]$ and let $\delta>0$.
+we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
 $\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2V_{n}(\bm{Z})}{n}\log\left(\frac{2}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{2}{\delta}\right),$
 where $V_{n}(\bm{Z}):=\frac{1}{n(n-1)}\sum_{1\leq i<j\leq n}(Z_{i}-Z_{j})^{2}$ is the sample variance.
-Lemma E.5 (Concentration of Chi-square variables).
-Let $Z_{1},\cdots,Z_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard normal. Then, it holds that
-${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|>\epsilon\right)\leq 2e^{-n\epsilon^{2}/8}.$
-Equivalently, with probability at least $1-\delta$, we have that
-$\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|\leq\sqrt{\frac{8}{n}\log\left(\frac{2}{\delta}\right)}$
-Lemma E.6 (Restatement of [43, Example 6.3]).
-Let $\bm{X}\in{\mathbb{R}}^{n\times d}$ be a random matrix from the $\Sigma$-Gaussian ensemble, i.e., each row vector $x_{i}$ is drawn i.i.d.
-from a multivariate ${\mathcal{N}}(0,\Sigma)$ distribution.
-Then, with probability at least $1-\delta$, we have that
-$\|\widehat{\Sigma}-\Sigma\|_{2}\leq\|\Sigma\|_{2}\left(2\sqrt{\frac{d}{n}}+2\epsilon+\left(\sqrt{\frac{d}{n}}+\epsilon\right)^{2}\right),\textup{ where }\epsilon=\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)},$ (36)
-where $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\top}$ is the sample covariance matrix.
-Lemma E.7 (Dimension-free concentration of sample covariance matrix [30, 25, 46]).
-Let $X_{1},\cdots,X_{n}$ are i.i.d zero-mean and sub-Gaussian random vectors with parameter at most $\sigma^{2}$, whose covariance matrix is given by $\Sigma$. Let
-$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (37)
-denote the effective rank. Then, with probability at least $1-\delta$, the sample covariance matrix satisfies that
-$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\sigma^{2}\|\Sigma\|_{2}\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}$ (38)
-whenever $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$.
-Lemma E.8.
-Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
-$\textup{Tr}(\sqrt{A+B})\geq\max\{\textup{Tr}(\sqrt{A}),\textup{Tr}(\sqrt{B})\}$ (39)
-Proof.
-Since $A+B\succeq A$ and $A+B\succeq B$, by Courant-Fischer min-max theorem, it holds that $\lambda_{i}(A+B)\geq\lambda_{i}(A)$ and $\lambda_{i}(A+B)\geq\lambda_{i}(B)$ for their $i$-th largest eigenvalues, which concludes the proof.
-∎
-Lemma E.9.
-Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
-$\textup{Tr}[\sqrt{A+B}]\leq\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}].$ (40)
-Proof.
-First, we assume that both matrices $A$ and $B$ are positive strictly definite. Then,
-$\textup{Tr}(\sqrt{A+B})=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}(A+B)(A+B)^{-\frac{1}{4}}]$
-$=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}A(A+B)^{-\frac{1}{4}}]+\textup{Tr}[(A+B)^{-\frac{1}{4}}B(A+B)^{-\frac{1}{4}}]$
-$=$ $\textup{Tr}[\sqrt{A}(A+B)^{-\frac{1}{2}}\sqrt{A}]+\textup{Tr}[\sqrt{B}(A+B)^{-\frac{1}{2}}\sqrt{B}]$
-$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
-where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma E.12). For general PSD matrices $A$ and $B$, note that
-$\textup{Tr}(\sqrt{A+B+\epsilon I})\leq\textup{Tr}(\sqrt{A+\epsilon I})+\textup{Tr}(\sqrt{B+\epsilon I})$
-holds for any $\epsilon>0$ by our previous analysis. Since the trace map is continuous, ∎
-Lemma E.10.
-Let $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
-$\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}\leq\sqrt{\frac{32}{n^{2}}\log\left(\frac{2}{\delta}\right)}.$
-Proof.
-Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$. By Lemma E.5, we have that
-${\mathbb{P}}\left(\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|>\epsilon\right)=$ ${\mathbb{P}}\left(|U^{2}-1|>n\epsilon\right)\leq 2\exp\left(-\frac{n^{2}\epsilon^{2}}{8}\right)$
-where $U^{2}\sim\chi^{2}_{1}$ is a chi-square random variable. Hence, with probability at least $1-\delta$, it holds that
-$\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|\leq\sqrt{\frac{8}{n^{2}}\log\left(\frac{2}{\delta}\right)},$
-which concludes the proof. ∎
-Lemma E.11.
-Let $z_{1},\cdots,z_{n},u_{1},\cdots,u_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
-$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|<\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)}.$
-Proof.
-Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$.
-$\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)=\frac{1}{n}\left(\frac{\sum_{i=1}^{n}z_{i}}{\sqrt{n}}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{\sqrt{n}}\right)=\frac{1}{n}ZU,$
-where $Z,U\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Observe that
-$ZU=\frac{1}{4}(Z+U)^{2}-\frac{1}{4}(Z-U)^{2}\sim\frac{1}{2}Q-\frac{1}{2}R,$
-where $Q:=\frac{1}{2}(Z+U)^{2},R:=\frac{1}{2}(Z-U)^{2}\sim\chi_{1}^{2}$ are i.i.d Chi-square distributed with 1 degree of freedom (since $\frac{1}{\sqrt{2}}(Z+U)$ and $\frac{1}{\sqrt{2}}(Z-U)$ are i.i.d standard Gaussian).
-Hence, by Lemma E.5 and a union bound, with probability at least $1-\delta$,
-$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|=$ $\left|\frac{1}{2}Q-\frac{1}{2}R\right|\leq\frac{1}{2}\left|Q-1\right|+\frac{1}{2}\left|R-1\right|$
-$\leq$ $\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)},$
-which concludes the proof. ∎
-Lemma E.12.
-Let $A,B$ denote two PSD matrices such that $A\succeq B$. Then, their inverse $B^{-1}\succeq A^{-1}$.
-Proof.
-It suffices to show that $\sqrt{A}B^{-1}\sqrt{A}\succeq I$.
-Since $\sqrt{A}B^{-1}\sqrt{A}=(\sqrt{A}B^{-\frac{1}{2}})(\sqrt{A}B^{-\frac{1}{2}})^{\top}$ shares the same eigenvalues with $(\sqrt{A}B^{-\frac{1}{2}})^{\top}(\sqrt{A}B^{-\frac{1}{2}})=B^{-\frac{1}{2}}AB^{-\frac{1}{2}}\succeq I$, which concludes the proof.
-∎
+Appendix C EXPERIMENTAL DETAILS
+1. List of pretrained generative models.
+For the CIFAR10 dataset, we compare pretrained generative models including iDDPM-DDIM (Nichol and Dhariwal,, 2021), LOGAN (Wu et al.,, 2020), WGAN-GP (Gulrajani et al.,, 2017), NVAE (Vahdat and Kautz,, 2020), and RESFLOW (Chen et al.,, 2019).
+For ImageNet, we compare pretrained models including DiT-XL-2 (Peebles and Xie,, 2023), ADMG (Dhariwal and Nichol,, 2021), BigGAN (Brock et al.,, 2019), RQ-Transformer (Lee et al.,, 2022), and ADM (Dhariwal and Nichol,, 2021).
+For the FFHQ dataset, we compare StyleNAT (Walton et al.,, 2023), StyleGAN2-ADA (Karras et al.,, 2020), LDM (Rombach et al.,, 2022), Unleashing-Transformers (Bond-Taylor et al.,, 2021), and Efficient-vdVAE (Hazami et al.,, 2022).
+We utilize the generated image datasets downloaded from the dgm-eval repository (Stein et al.,, 2023).
+2. List of embeddings for FD-based evaluation and selection.
+We consider three standard encoders for image data: InceptionV3.Net (Szegedy et al.,, 2016), DINOv2 (Oquab et al.,, 2024), and CLIP (Radford et al.,, 2021).
+Following (Stein et al.,, 2023), we utilize the DINOv2 ViT-L/14 model and the OpenCLIP ViT-L/14 trained on DataComp-1B (Gadre et al.,, 2023).
+Embeddings extracted by InceptionV3.Net, DINOv2, and CLIP have $2048$, $1024$, and $1024$ dimensions, respectively.
+For video data, we utilize the I3D model (Carreira and Zisserman,, 2018) pretrained on the Kinetics-400 (Carreira and Zisserman,, 2018) dataset following (Unterthiner et al.,, 2019), where the logits layer with $400$ dimensions is used as the embedding.
+For audio data, we utilize VGGish (Hershey et al.,, 2017) following (Kilgour et al.,, 2019), where the activations from the $128$ dimensional layer prior to the final classification layer are used as the embedding.
+3. Implementation details of FD-UCB.
+As we consider InceptionV3, DINOv2, and (unnormalized) CLIP embeddings, which are generally unbounded, we utilize the collected data to estimate the model-dependent parameters in the bonus (6).
+This approach preserves the online format of the FD-UCB algorithm.
+Our numerical results indicate that with only 50 samples, the norm terms in the bound would be close to their underlying values (Tables 12, and 3).
+To have a better estimation of the covariance matrix, we also adopt the thresholding method in (Cai and Liu,, 2011).
+We treat the batch size, parameter $\kappa$ in the bonus (6), and parameter $M$ for thresholding as hyperparameters.
+We conduct ablation study on these hyperparameters summarize the results in Figure 14.
+Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
+DiT-XL-2 $166.4$ $160.0$ $12.5$ $20.9$ $9.6$ $11.2$ $17.3$ $14.3$ $[9.7,33.6]$
+ADMG $157.7$ $150.8$ $12.1$ $19.4$ $9.2$ $10.5$ $17.2$ $14.4$ $[10.8,34.5]$
+BigGAN $154.5$ $161.8$ $13.5$ $22.3$ $11.0$ $13.2$ $14.0$ $12.2$ $[9.7,31.8]$
+RQ-Transformer $166.9$ $164.5$ $13.8$ $22.2$ $10.7$ $12.7$ $15.5$ $13.0$ $[11.0,29.7]$
+ADM $176.5$ $182.5$ $15.7$ $25.6$ $12.9$ $15.2$ $13.7$ $12.0$ $[0.5,34.8]$
+ImageNet $181.2$ $141.2$ $12.4$ $20.2$ $9.2$ $8.7$ $19.7$ $16.2$ $[10.9,33.0]$
+Table 1: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively.
+The image data embeddings are extracted by InceptionV3.Net.
+Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
+DiT-XL-2 $591.2$ $601.9$ $59.1$ $101.4$ $38.2$ $48.0$ $15.5$ $12.5$ $[34.6,36.9]$
+ADMG $579.9$ $566.1$ $62.1$ $101.3$ $41.2$ $50.4$ $14.1$ $11.2$ $[34.9,37.0]$
+BigGAN $514.9$ $486.8$ $62.7$ $90.8$ $43.8$ $54.0$ $11.8$ $9.0$ $[34.9,36.8]$
+RQ-Transformer $545.5$ $551.7$ $58.8$ $99.1$ $39.3$ $51.0$ $13.9$ $10.8$ $[35.0,36.8]$
+ADM $560.2$ $561.8$ $59.0$ $100.7$ $38.3$ $52.5$ $14.6$ $10.7$ $[34.9,36.8]$
+ImageNet $636.8$ $573.1$ $53.5$ $118.5$ $32.1$ $74.1$ $19.9$ $7.7$ $[35.1,36.7]$
+Table 2: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively.
+The image data embeddings are extracted by CLIP.
+Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
+DiT-XL-2 $2079.5$ $2085.6$ $99.9$ $309.4$ $36.5$ $81.4$ $57.0$ $25.6$ $[38.5,48.7]$
+ADMG $2068.2$ $2102.1$ $103.0$ $311.4$ $42.5$ $79.1$ $48.7$ $25.6$ $[38.6,49.0]$
+BigGAN $1885.5$ $1799.6$ $120.3$ $285.4$ $68.8$ $110.4$ $27.4$ $16.3$ $[36.5,48.6]$
+RQ-Transformer $1955.3$ $1968.0$ $111.4$ $302.5$ $54.4$ $91.1$ $36.0$ $21.6$ $[38.4,49.3]$
+ADM $2013.3$ $1988.1$ $106.7$ $306.0$ $50.5$ $101.0$ $39.9$ $19.7$ $[39.1,49.2]$
+ImageNet $2112.7$ $2095.4$ $85.3$ $347.3$ $21.0$ $104.3$ $100.7$ $20.1$ $[42.2,48.3]$
+Table 3: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively.
+The image data embeddings are extracted by DINOv2.
+4. Implementation details of Naive-UCB.
+Naive-UCB is a simplification of FD-UCB and IS-UCB which replaces the generator-dependent variables in the bonus function with data-independent and dimension-based terms.
+For FD-based evaluation, Naive-UCB sets $\textup{Tr}[\Sigma_{g}]=O(d),\textup{Tr}[\Sigma_{g}^{2}]=O(d)$, and $\|\Sigma_{g}\|_{2}=O(1)$ in FD-UCB.
+For IS-based evaluation, the Naive-UCB method sets $\widehat{V}^{n}(Y_{g}|X_{g})=(\log d)^{2}$ and $\widehat{V}^{n}(p_{Y|X_{g}}[j])=1$ for any $j\in[d]$.
+Avg. Regret
+Figure 7: Online FD-based evaluation and selection among standard generative models: The image data embeddings are extracted by InceptionV3.Net. Results are averaged over 20 trials.
+[…]
+Avg. Regret
+Figure 8: Online FD-based evaluation and selection among standard generative models: The image data embeddings are extracted by DINOv2-ViT-L/14. Results are averaged over 20 trials.
+[…]
+FFHQ (VC)
+[…]
+Figure 9: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by InceptionV3.Net. Results are averaged over 20 trials.
+FFHQ (VC)
+[…]
+Figure 10: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by DINOv2. Results are averaged over 20 trials.
+FFHQ (VC)
+[…]
+Figure 11: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by CLIP. Results are averaged over 20 trials.
+MSR-VTT (Video)
+[…]
+Figure 12: Online evaluation and selection on video and audio data: The video data embeddings are extracted by I3D, and the audio data embeddings are extracted by VGGish.
+Results are averaged over 20 trials.
+Figure 13: Online IS-based evaluation and selection among variance-controlled models on the AFHQ Dog dataset: IS-UCB can identify models that generate images with more diversity.
+Results are averaged over 20 trials.
+Batch Size
+Figure 14: Ablation study on the hyperparameters of FD-UCB on the CIFAR10 dataset: The image data embeddings are extracted by CLIP. Results are averaged over 20 trials.
+Parameter $\kappa$
+[…]
+Magnatagatune (Audio)
+[…]
+Figure 15: Adaptation to new models: A new model is introduced after each $200$ steps. Results are averaged over 20 trials.